### 10 Mar 2026 v1

# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

Abstract
Model merging has emerged as a transformative paradigm for combining the capabilities of multiple neural networks into a single unified model without additional training. With the rapid proliferation of fine-tuned large language models (LLMs), merging techniques offer a computationally efficient alternative to ensembles and full retraining, enabling practitioners to compose specialized capabilities at minimal cost. This survey presents a comprehensive and structured examination of model merging in the LLM era through the FUSE taxonomy, a four-dimensional framework organized along Foundations, Unification Strategies, Scenarios, and Ecosystem. We first establish the theoretical underpinnings of merging, including loss landscape geometry, mode connectivity, and the linear mode connectivity hypothesis. We then systematically review the algorithmic landscape, spanning weight averaging, task vector arithmetic, sparsification-enhanced methods, mixture-of-experts architectures, and evolutionary optimization approaches. For each method family, we analyze the core formulation, highlight representative works, and discuss practical trade-offs. We further examine downstream applications across multi-task learning, safety alignment, domain specialization, multilingual transfer, and federated learning. Finally, we survey the supporting ecosystem of open-source tools, community platforms, and evaluation benchmarks, and identify key open challenges including theoretical gaps, scalability barriers, and standardization needs. This survey aims to equip researchers and practitioners with a structured foundation for advancing model merging.
1 Introduction
Model merging combines the parameters of multiple trained neural networks into a single unified model, inheriting their capabilities without inference-time ensemble overhead (Li et al., 2023a). Unlike ensembles that aggregate predictions from separate models at runtime, merging operates directly in weight space, yielding a single model whose cost matches any individual source. Early work showed that averaging parameters of models trained with different random seeds improves generalization performance (Izmailov et al., 2018). In the early days, these techniques were primarily applied to homogeneous models trained on identical datasets, with Stochastic Weight Averaging emerging as a principled approach for combining checkpoints along a single optimization trajectory (Garipov et al., 2018). With the advent of large-scale pretrained language model (Devlin et al., 2019)s, the scope of model merging expanded dramatically, enabling the combination of models fine-tuned for entirely different tasks into unified systems exhibiting multi-task capabilities (Ilharco et al., 2023). This evolution from checkpoint averaging to cross-task composition marks a shift in how the community conceptualizes neural network knowledge reuse.
The emergence of pretrained foundation models such as LLaMA (Touvron et al., 2023a), Mistral (Jiang et al., 2023), Qwen (Hu et al., 2023), DeepSeek (Bi et al., 2024), Gemma (Team, 2024a), and CLIP (Radford et al., 2021) has created conditions that make model merging both practically viable and scientifically compelling. The key insight is that models fine-tuned from a shared pretrained initialization tend to reside within the same loss basin, exhibiting linear mode connectivity (Singh and Jaggi, 2020) that permits direct weight interpolation without traversing high-loss barriers (Garipov et al., 2018). This theoretical grounding, combined with the proliferation of openly available fine-tuned large language models, has sparked substantial interest in merging methodologies. The past two years have witnessed an explosion of research activity, with novel techniques such as task vector arithmetic, sparsification-enhanced merging, and activation-informed (Hammoud et al., 2024) approaches achieving notable success in combining specialized models (Yadav et al., 2023; Yu et al., 2023; Huang et al., 2024). Practically, merged models have achieved top rankings on competitive leaderboards including the Open LLM Leaderboard (Li et al., 2023b), demonstrating that strategic model combination can yield capabilities exceeding those of individual fine-tuned variants (Wortsman et al., 2022a; Jang et al., 2024). The establishment of dedicated toolkits such as mergekit (Goddard et al., 2024) has further democratized access to these techniques, enabling experiment with sophisticated merging strategies without extensive technical expertise. This confluence of theoretical understanding, empirical success, and practical tooling (Goddard et al., 2024) has positioned model merging as a central technique in modern LLM development.
Several recent surveys have begun to address aspects of this rapidly evolving landscape. Yang et al. (2024b) and Li et al. (2023a) provide a broad treatment of model fusion techniques across deep learning, while Yadav et al. (2024a) focus specifically on mixture-of-experts style composition approaches. Yang et al. (2024b) examine collaborative strategies more broadly, encompassing both merging and ensemble methods. However, existing surveys lack a unified treatment that systematically connects theoretical foundations explaining why merging succeeds, algorithmic strategies for performing merging effectively, downstream application scenarios where merging provides value, and the practical ecosystem supporting deployment. No prior work has comprehensively addressed the full spectrum from loss landscape geometry and mode connectivity theory through to evaluation methodologies and community tooling, nor proposed a unified taxonomy for navigating this space.
This survey provides a systematic treatment of model merging in the LLM era, bridging theoretical understanding with practical application. Table 1 summarizes the major open-weight foundation models (Brown et al., 2020; Team, 2024b) commonly used in merging research. As shown in Figure 2, our survey is organized around the FUSE taxonomy, a four-dimensional framework encompassing Foundations, Unification Strategies, Scenarios, and Ecosystem, that structures the field according to why merging works, how merging is performed, where merging provides value, and what resources support merging in practice. Figure 1 provides a schematic overview of the model merging pipeline. The main contributions of this survey are as follows: (1) We propose the FUSE taxonomy as a principled organizational framework that comprehensively categorizes model merging research across theoretical foundations including loss landscape geometry, mode connectivity, and weight space symmetries; algorithmic unification strategies spanning weight-space averaging, task vector arithmetic, geometric interpolation, and architecture-level composition; application scenarios including capability augmentation, alignment and safety, and efficiency-driven deployment; and ecosystem infrastructure encompassing toolkits, benchmark s, and community resources. (2) We provide an in-depth technical analysis of merging methodologies, elucidating the mathematical principles underlying each approach and offering comparative insights into their respective strengths, limitations, and appropriate use cases. (3) We conduct a systematic examination of application domains, synthesizing empirical evidence regarding when and how model merging delivers practical benefits across multi-task generalization, multilingual transfer, preference alignment, and federated learning scenarios. (4) We identify open challenges and articulate promising future research (Li et al., 2025) directions, including theoretical gaps in understanding mergeability, scalability limitations for increasingly large models, and opportunities for automated and dynamic merging systems. The remainder of this survey proceeds as follows: Section 2 establishes theoretical foundations; Sections 3–5 present unification strategies with increasing sophistication; Section 6 examines application scenarios; and Section 7 discusses the ecosystem, open challenges, and future directions.
Table 1: Overview of major open-weight foundation models (Brown et al., 2020; Team, 2024b) commonly used in model merging research. These models serve as the base for fine-tuning and subsequent merging experiments.
Model Family Organization Parameters Year Architecture Key Features
LLaMA (Touvron et al., 2023a) Meta 7B–65B 2023 Decoder-only RoPE, Pre-norm, SwiGLU
LLaMA-2 (Touvron et al., 2023b) Meta 7B–70B 2023 Decoder-only GQA, Extended context
LLaMA-3 (Team, 2024b) Meta 8B–405B 2024 Decoder-only 128K context, GQA
Mistral (Jiang et al., 2023) Mistral AI 7B 2023 Decoder-only Sliding window attention
[…]
Qwen (Hu et al., 2023) Alibaba 1.8B–72B 2023 Decoder-only Multilingual, Code
Qwen2 (Yang et al., 2024a) Alibaba 0.5B–72B 2024 Decoder-only GQA, Extended context
DeepSeek (Bi et al., 2024) DeepSeek 7B–67B 2024 Decoder-only MHA, Scaling law design
DeepSeek-V2 (DeepSeek-AI, 2024) DeepSeek 236B (21B active) 2024 MoE MLA, DeepSeekMoE
Gemma (Team, 2024a) Google 2B–7B 2024 Decoder-only RoPE, Multi-Query Attn
[…]
Figure 1: Overview of the model merging pipeline. Task-specific models $\{\theta_{1},\dots,\theta_{K}\}$, obtained by fine-tuning a shared pretrained checkpoint $\theta_{0}$, are combined via one of several strategies (Sections 3–5) into a single model $\theta_{\mathrm{merged}}$, which is then evaluated on all source tasks. Gray vertical bars represent the many-to-many mapping between model inputs and merging strategies.
2 Theoretical Foundations of Model Merging
FUSE Taxonomy for Model MergingF – Foundations • Loss Landscape • Mode Connectivity • Perm. Symmetry • Prerequisites U – Unification • Weight Averaging • Task Vectors • MoE Merging • Search-Based S – Scenarios • Multi-Task • Safety & Alignment • Federated Learning • Continual Adapt. E – Ecosystem • Toolkits • Benchmarks • Future Directions
Figure 2: The FUSE taxonomy. Foundations (§2): why merging works. Unification (§3–§5): how models are combined. Scenarios (§6): where merging applies. Ecosystem (§7): what supports deployment.
Why does model merging work? We examine three interconnected pillars: the geometric structure of loss landscapes that permits effective weight interpolation (Ren et al., 2025), the phenomenon of mode connectivity enabling low-loss paths between independently trained solutions (Draxler et al., 2018; Lubana et al., 2023), and weight-space symmetries, particularly permutation invariance (Jordan et al., 2022; Entezari et al., 2022), that must be resolved for effective merging (Ainsworth et al., 2023).
2.1 Loss Landscape Geometry and Convexity Properties in Neural Networks
The success of model merging fundamentally depends on the geometric properties of the loss landscape in which neural network parameters reside. Understanding this landscape, particularly how loss values vary as one traverses the high-dimensional parameter space, provides essential insight into why weight interpolation between independently trained models can yield functional merged networks rather than catastrophic performance degradation.
Neural network loss landscapes exhibit remarkably complex topologies characterized by numerous local minima, saddle points, and regions of varying curvature. However, a crucial observation that underpins model merging is that modern deep networks, particularly when sufficiently overparameterized, display favorable geometric properties that deviate significantly from the pathological non-convex landscapes one might naively expect. Singh et al. (2024) provide a systematization of these landscape properties, demonstrating that overparameterization fundamentally alters the geometry of the optimization surface. Specifically, when parameters substantially outnumber the minimum required for the task, the loss landscape exhibits large connected regions of near-optimal solutions rather than isolated minima.
The concept of basin characteristics, encompassing both width and flatness, plays a central role in determining whether weight interpolation between two solutions will preserve model functionality. A loss basin can be formally characterized by considering the set of parameters $\mathcal{B}_{\epsilon}(\theta^{*})=\{\theta:\mathcal{L}(\theta)<\mathcal{L}(\theta^{*})+\epsilon\}$ surrounding a local minimum $\theta^{*}$. The width of this basin, often measured through the eigenspectrum of the Hessian matrix $\nabla^{2}\mathcal{L}(\theta^{*})$, determines how much perturbations from the optimum increase loss. Wide, flat basins permit substantial movement in parameter space without significant performance degradation, directly enabling weight averaging.
Garipov et al. (2018) investigated the empirical phenomenon of monotonic linear interpolation, wherein the loss along the linear path $\theta_{\alpha}=(1-\alpha)\theta_{1}+\alpha\theta_{2}$ between two trained networks exhibits monotonically decreasing behavior toward the interpolated midpoint under certain conditions. This monotonicity property suggests that linear interpolation not only avoids catastrophic loss barriers but can actually discover solutions with lower loss than either endpoint, a finding with important implications for model merging strategies that rely on simple averaging operations.
The relationship between loss landscape geometry and generalization performance introduces additional considerations for model merging. Flat minima, characterized by low curvature in the Hessian eigenspectrum, have been associated with improved generalization properties  (Izmailov et al., 2018). When merging models that reside in flat regions of the loss landscape, the resulting averaged parameters are more likely to remain within a low-loss region, thereby preserving generalization capabilities. Conversely, sharp minima (Keskar et al., 2017) with high curvature present challenges for merging, as even small perturbations induced by averaging can push parameters into high-loss regions.
Lubana et al. (2023) further examined the dynamics of weight interpolation, revealing that interpolated models can exhibit transient performance characteristics that differ from simple averaging predictions. Their analysis demonstrates that the geometry of the loss landscape along interpolation paths depends critically on training dynamics and the specific optimization trajectory taken to reach each solution.
Mathematical analysis of overparameterized networks reveals that the loss surface increasingly resembles a high-dimensional valley structure as model capacity grows. Let $\mathcal{M}$ denote the manifold of global minima; in the overparameterized regime, $\mathcal{M}$ becomes a connected, often convex subspace (Tang et al., 2024d) within which arbitrary linear combinations of solutions remain optimal. This geometric insight, explored in (Ferbach et al., 2024), provides theoretical justification for why weight averaging succeeds when models share sufficient structural commonality, their solutions lie within the same connected low-loss region, rendering interpolation a geometrically valid operation (Neyshabur et al., 2020).
2.2 Linear Mode Connectivity Theory and Basin Characteristics
The geometric insight that successful merging requires solutions to occupy connected low-loss regions naturally leads to the formal framework of mode connectivity, which provides a fundamental theoretical lens through which to understand when and why model merging succeeds. Mode connectivity refers to the existence of paths in weight space connecting different trained solutions along which the loss remains consistently low. This concept has emerged as central to explaining the empirical success of weight interpolation methods, as thoroughly surveyed in (Yang et al., 2024b).
Formally, two neural network solutions $\theta_{1}$ and $\theta_{2}$ are said to exhibit linear mode connectivity (Singh and Jaggi, 2020) if the loss along the linear interpolation path $\theta(\alpha)=(1-\alpha)\theta_{1}+\alpha\theta_{2}$ for $\alpha\in[0,1]$ satisfies $\mathcal{L}(\theta(\alpha))\leq\max(\mathcal{L}(\theta_{1}),\mathcal{L}(\theta_{2}))+\epsilon$ for some small tolerance $\epsilon\geq 0$. When $\epsilon=0$, the models reside in the same loss basin with no interpolation barrier. More generally, nonlinear mode connectivity (Singh and Jaggi, 2020) relaxes the path constraint to allow curved trajectories, typically parameterized as Bézier curves or piecewise linear paths. This distinction has practical implications: linear mode connectivity (Singh and Jaggi, 2020) enables direct weight averaging, while nonlinear connectivity may require more sophisticated merging procedures.
The analysis of loss barriers, the maximum loss increase encountered along interpolation paths, provides quantitative measures for characterizing mode connectivity strength. Singh et al. (2024) present a systematization of methods for computing and interpreting these barriers, noting that barrier height correlates inversely with merging success. Specifically, given solutions $\theta_{1}$ and $\theta_{2}$, the loss barrier is defined as:
$\Delta\mathcal{L}=\max_{\alpha\in[0,1]}\mathcal{L}(\theta(\alpha))-\frac{1}{2}(\mathcal{L}(\theta_{1})+\mathcal{L}(\theta_{2}))$ (1)
When $\Delta\mathcal{L}\leq\epsilon$ for some small threshold $\epsilon$, we say the models exhibit $\epsilon$-linear mode connectivity (Singh and Jaggi, 2020). The relationship between loss barrier magnitude and merging success can be formally characterized through the following theorem:
Proposition 1 (Merging Error Bound (Singh et al., 2024)).
Let $\theta_{1},\theta_{2}$ be two models with shared pretrained initialization $\theta_{0}$, and let $\theta_{\text{avg}}=\frac{1}{2}(\theta_{1}+\theta_{2})$. Under $L$-smoothness of the loss function $\mathcal{L}$, the performance degradation of the averaged model is bounded by:
$\mathcal{L}(\theta_{\text{avg}})-\frac{1}{2}(\mathcal{L}(\theta_{1})+\mathcal{L}(\theta_{2}))\leq\frac{L}{8}\|\theta_{1}-\theta_{2}\|^{2}$ (2)
This bound reveals why shared initialization is critical: models fine-tuned from the same pretrained checkpoint exhibit smaller $\|\theta_{1}-\theta_{2}\|$, yielding tighter merging guarantees.
A critical insight enabling practical model merging in the large language model era concerns the role of shared pretrained initialization. When multiple models are fine-tuned from a common pretrained checkpoint, they exhibit substantially stronger linear mode connectivity (Singh and Jaggi, 2020) compared to independently trained models (Frankle et al., 2020; Garipov et al., 2018). This phenomenon arises because fine-tuning typically induces relatively small parameter perturbations that remain within the same loss basin as the pretrained model, whereas training from random initialization explores fundamentally different regions of the loss landscape.
Empirical investigations have employed various path parameterizations to probe connectivity structure. Quadratic Bézier curves, parameterized as $\theta(\alpha)=(1-\alpha)^{2}\theta_{1}+2\alpha(1-\alpha)\theta_{m}+\alpha^{2}\theta_{2}$ with learnable midpoint $\theta_{m}$, can discover low-loss nonlinear paths even when linear barriers exist. Ren et al. (2025) demonstrated that trained networks frequently exhibit monotonic interpolation properties, where loss decreases steadily from one endpoint toward the other, suggesting asymmetric basin structures.
The work of Nobari et al. (2025) revealed that apparent loss barriers often arise from feature distribution mismatches rather than genuine geometric separation, proposing activation renormalization to restore connectivity. This finding suggests that many observed failures of linear mode connectivity (Singh and Jaggi, 2020) stem from correctable statistical artifacts rather than fundamental incompatibility between solutions.
Understanding basin characteristics extends beyond barrier analysis to encompass basin width and shape. Lubana et al. (2023) demonstrated that interpolated models can exhibit performance dynamics reflecting the local basin geometry, with wider basins supporting more robust interpolation. These theoretical insights directly inform practical merging strategies by establishing that shared pretraining serves as a powerful mechanism for ensuring models occupy connected regions amenable to linear combination. However, even when models theoretically reside in connected basins, the success of direct weight averaging depends critically on how the models are represented, a consideration that leads naturally to examining the fundamental symmetries inherent in neural network weight spaces.
2.3 Weight Space Symmetries and Permutation Invariance
These fundamental symmetries, particularly permutation invariance, substantially impact the feasibility and methodology of model merging. Permutation invariance arises from that reordering hidden units within a layer, along with corresponding adjustments to incoming and outgoing weight matrices, produces a functionally equivalent network. This mathematical property implies that the number of distinct parameterizations encoding identical input-output mappings grows combinatorially with network depth and width, creating substantial challenges for direct weight space operations.
Consider a feedforward network with weight matrices $W^{(l)}$ connecting layer l1l-1 to layer $l$. For any permutation matrix $P^{(l)}$ applied to the hidden units of layer $l$, the transformation $(W^{(l)},W^{(l+1)})\mapsto(P^{(l)}W^{(l)},W^{(l+1)}(P^{(l)})^{T})$ preserves the network’s function exactly. For a network with $L$ hidden layers each containing $n$ units, this yields $(n!)^{L}$ equivalent parameterizations, a symmetry group whose size renders the loss landscape highly non-convex despite potential functional simplicity. The comprehensive treatment in (Ainsworth et al., 2023) emphasizes that this symmetry significantly complicates the interpretation of distances and interpolations in weight space, as two networks computing identical functions may appear arbitrarily distant under Euclidean metrics.
The implications for model merging are immediate and severe. When averaging weights from two independently trained networks, the hidden unit orderings are essentially arbitrary and unrelated, causing the averaged network to combine misaligned features in semantically meaningless ways. Formally, if $\theta_{A}$ and $\theta_{B}$ represent two solutions related by permutation such that $\theta_{B}=\pi(\theta_{A})$ for some permutation $\pi$, then $\frac{1}{2}(\theta_{A}+\theta_{B})$ generally lies far from both solutions in functional space despite occupying their geometric midpoint. This observation explains why naive weight averaging of independently trained networks typically produces catastrophic performance degradation, even when both source networks achieve comparable accuracy.
The theoretical framework connecting permutation alignment (Ainsworth et al., 2023) to optimal transport provides principled methods for resolving this challenge. Given two networks with weight matrices $W_{A}$ and $W_{B}$, finding the optimal permutation reduces to solving an assignment problem that minimizes some distance metric between aligned representations. Ferbach et al. (2024) discuss how this formulation enables tractable algorithms based on the Hungarian method or approximate transport solvers, establishing correspondence between hidden units based on their functional roles rather than arbitrary indexing.
The work of Nobari et al. (2025) demonstrated that even after permutation alignment, residual distribution mismatches in intermediate activations can create apparent barriers to interpolation. Their analysis revealed that permutation-aligned networks may still exhibit incompatible feature statistics, means and variances of hidden activations, that cause interpolated networks to produce out-of-distribution internal representations. This insight motivates post-alignment normalization procedures that reconcile activation distributions, further reducing barriers to successful merging.
Shared initialization provides a natural mechanism for symmetry breaking that circumvents permutation ambiguity entirely. When multiple models are fine-tuned from identical pretrained weights, the optimization trajectories implicitly preserve hidden unit correspondence established during pretraining. This explains the empirical observation that fine-tuned models exhibit substantially better mergeability than independently trained counterparts (Wortsman et al., 2022a). The pretrained initialization effectively anchors each hidden unit to a specific functional role, ensuring that corresponding parameters across fine-tuned variants encode semantically related modifications.
Recent systematic investigations have quantified the relationship between alignment quality and merging success, revealing that architectural choices and training procedures influence the degree to which permutation alignment is necessary. Transformer architectures with their structured attention mechanisms exhibit different symmetry properties than fully-connected networks. These insights regarding weight space symmetries, combined with the earlier discussions of loss landscape geometry and mode connectivity, collectively inform a unified understanding of the conditions necessary for successful model merging.
2.4 Prerequisites and Conditions for Successful Model Merging
This unified understanding can be formalized as a framework of mergeability conditions, the set of prerequisites that must be satisfied for model merging to yield high-quality results. Understanding these conditions is essential for practitioners seeking to predict whether a given set of models can be successfully combined and for researchers developing new merging algorithms with broader applicability.
The most critical prerequisite for successful model merging is shared pretrained initialization, which establishes a common reference point in weight space from which all candidate models diverge during fine-tuning. Formally, consider a pretrained model with parameters $\theta_{\text{pre}}$ and a set of fine-tuned models $\{\theta_{1},\theta_{2},\ldots,\theta_{K}\}$ each obtained by optimizing task-specific objectives starting from $\theta_{\text{pre}}$. The shared initialization ensures that fine-tuned parameters remain within the same loss basin, enabling linear interpolation $\theta_{\alpha}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}$ to traverse low-loss regions rather than crossing high-loss barriers (Garipov et al., 2018; Frankle et al., 2020). This basin-sharing property significantly distinguishes the fine-tuning paradigm from independent training, where models converge to disconnected minima separated by substantial energy barriers.
Architectural compatibility is a second essential condition, requiring both structural identity and representational alignment. Models must share identical architectures—layer configurations, activation functions, and normalization schemes—for element-wise parameter operations to be well-defined. Beyond this structural requirement, Li et al. (2023a) demonstrate that certain architectural choices exhibit inherently favorable merging properties. Transformer architectures (Vaswani et al., 2017), with their residual connections and layer normalization, maintain more consistent feature representations across fine-tuning variants, facilitating parameter combination without the severe interference observed in earlier network designs. The residual stream structure effectively constrains the magnitude of per-layer modifications, keeping fine-tuned models closer to the pretrained initialization.
Training procedure similarity also strongly influences merging success. Models fine-tuned with similar hyperparameters (learning rate, batch size, optimizer) tend to follow comparable trajectories through weight space, yielding stronger linear mode connectivity (Singh and Jaggi, 2020). Conversely, substantial procedural divergence can cause models to settle in distinct sub-basins within the broader pretrained basin, introducing interpolation barriers that degrade merge quality. Izmailov et al. (2018) provide empirical analysis showing how training duration and learning rate magnitude influence the distance between fine-tuned solutions and consequently the loss barrier encountered during interpolation.
The theoretical framework for predicting capability preservation during merging remains incomplete but can be partially characterized through the lens of task vector interference. When fine-tuning induces parameter modifications in complementary subspaces, different parameters or opposing directions for different tasks, merging preserves individual capabilities through linear superposition. However, when task-specific adaptations overlap substantially and conflict in direction, the merged model exhibits degraded performance on all constituent tasks. Nobari et al. (2025) demonstrate that even when parameter-level conflicts are resolved, activation distribution mismatches can cause capability degradation, highlighting the multi-faceted nature of mergeability assessment. This observation motivates the development of mergeability prediction methods that analyze both parameter-space and representation-space compatibility, enabling principled decisions about which models can be successfully combined. Yet while these conditions provide practical guidance, the theoretical foundations underlying them, particularly for large-scale language models, remain incompletely understood.
2.5 Open Theoretical Questions and Research Frontiers
Despite this progress, several substantial gaps persist between observed phenomena and rigorous theoretical explanations, presenting fertile ground for future research. The most pressing open question concerns the fundamental mechanisms by which pretrained large language models exhibit such strong linear mode connectivity (Singh and Jaggi, 2020) following independent fine-tuning on disparate tasks. While it is well-established that shared initialization creates a common loss basin, the theoretical understanding of why this basin remains sufficiently broad and convex to accommodate diverse task-specific adaptations, particularly as model scale increases to billions of parameters, remains incomplete. (Singh et al., 2024) identifies this explanatory gap as a central challenge, noting that existing mode connectivity theory was largely developed in smaller neural networks and may require substantial extension to account for the unique properties of large-scale language models.
The relationship between model scale and mergeability properties presents another theoretically unresolved frontier. Empirical observations suggest that larger models exhibit more favorable merging characteristics, with reduced loss barriers and improved capability retention compared to smaller counterparts. However, a principled mathematical framework explaining this phenomenon remains elusive. One hypothesis posits that increased overparameterization expands the effective dimensionality of the loss basin, providing more ”room” for task-specific adaptations to coexist without interference. An alternative perspective suggests that the emergent representations in larger models may be more compositional and modular, naturally supporting the superposition of task-specific modifications. Recent work provides empirical investigation into scale effects but acknowledges that theoretical explanations for observed scaling behaviors remain speculative.
Emerging research directions seek to develop predictive frameworks for merge compatibility that could obviate the need for exhaustive experimentation. The challenge can be formalized as learning a function $\mathcal{C}:(\theta_{1},\theta_{2},\ldots,\theta_{k})\rightarrow[0,1]$ that estimates the expected performance of merging models $\theta_{1},\ldots,\theta_{k}$ without actually performing the merge. Promising approaches analyze task vector statistics, including angular relationships and magnitude distributions, to predict interference severity. surveys existing techniques for characterizing solution relationships through loss landscape analysis, providing foundational tools that could support compatibility prediction. However, translating these geometric insights into reliable predictive models requires addressing the computational challenge of efficiently approximating high-dimensional loss surface properties.
Theoretical investigations into the fundamental limits of merging constitute another critical research frontier. While sparsification-based methods such as TIES-Merging and DARE have demonstrated that parameter interference can be mitigated, the question of when interference becomes irreconcilable, representing a fundamental rather than technical limitation, remains open. Nobari et al. (2025) demonstrate that activation-level statistics provide complementary information to parameter-level analysis, suggesting that comprehensive mergeability theory must account for both weight space and representation space phenomena. Furthermore, the conditions under which nonlinear interpolation paths offer strictly superior outcomes compared to linear combinations, and how such paths might be efficiently discovered, represent underexplored theoretical territories with significant practical implications.
Finally, extending merging theory to accommodate architectural heterogeneity presents significant theoretical challenges. Current frameworks assume exact architectural correspondence, but practical scenarios increasingly involve models with variations in layer counts, hidden dimensions, or attention mechanisms. Developing principled theoretical foundations for cross-architecture merging, potentially leveraging representation alignment or functional correspondence rather than direct parameter mapping, could substantially expand the applicability of model merging techniques while requiring novel mathematical frameworks that transcend current weight-space formulations.
Table 2: Comparison of theoretical foundations and concepts in model merging.
Theoretical Concept Key Approach Key Feature(s) Limitations/Notes Year
Loss Landscape Geometry Analyzes geometric properties of high-dimensional parameter space Explains why weight interpolation yields functional networks Complex landscapes make complete characterization difficult 2024
Linear Mode Connectivity
(Garipov et al., 2018)
Identifies low-loss paths between trained solutions in weight space Central to explaining success of weight interpolation methods Requires shared initialization for strong connectivity 2024
Permutation Invariance
(Ainsworth et al., 2023)
Exploits symmetry from reordering hidden units with weight adjustments Multiple equivalent parameterizations for same function Alignment required before merging independently trained models 2024
Weight Space Symmetries
(Ainsworth et al., 2023)
Analyzes fundamental symmetry properties affecting merging feasibility Enables functionally equivalent network representations Symmetry breaking needed for meaningful comparisons 2024
Mergeability Conditions
(Singh et al., 2024)
Establishes unified framework of prerequisites for successful merging Predictive criteria for merge quality assessment Conditions may vary across model architectures and tasks 2024
Shared Initialization Theory
(Singh and Jaggi, 2020)
Explains role of common pretrained weights in enabling merging Strong predictor of linear mode connectivity Theoretical mechanisms not fully understood for LLMs 2024
Convexity Analysis Examines local and global convexity properties of loss surfaces Convex regions enable reliable weight averaging Neural networks generally non-convex globally 2024
The theoretical foundations examined in this section, encompassing loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and weight space symmetries, collectively establish the mathematical conditions under which model merging can succeed (summarized in Table 2), while also highlighting the fundamental barriers that arise when models reside in disconnected basins or exhibit misaligned permutation symmetries. These insights provide essential guidance for understanding when simple averaging will preserve model capabilities and when more sophisticated interventions become necessary. With this theoretical groundwork established, the following section turns to practical weight-space averaging and geometric interpolation methods, which translate these foundational principles into the algorithmic methods that instantiate these principles.
3 Weight-Space Averaging and Geometric Interpolation Methods
This section examines weight-space averaging and geometric interpolation—the most fundamental class of merging techniques that operate directly on model parameters. We organize these approaches along a spectrum of geometric sophistication: static averaging with fixed coefficients, trajectory-based methods that exploit optimization dynamics, and geometric interpolation (e.g., SLERP) that respects the manifold structure of weight spaces.
3.1 Linear Averaging and Static Model Soups
Weight averaging represents the most fundamental instantiation of model merging. It computes the element-wise arithmetic mean of model parameters, treating weight space as a Euclidean vector space. This formula admits a natural generalization to weighted averaging $\theta_{\text{avg}}=\sum_{i}\alpha_{i}\theta_{i}$ with $\sum_{i}\alpha_{i}=1$, where the coefficients $\alpha_{i}$ can be tuned on a held-out set or determined via importance metrics such as Fisher information (Matena and Raffel, 2022). Given a collection of $N$ models with parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$, uniform weight averaging produces a merged model with parameters:
$\theta_{\text{merged}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$
This formulation assumes all source models share identical architectures with correspondingly positioned parameters. This works because of linear mode connectivity (Singh and Jaggi, 2020): models fine-tuned from a common pretrained initialization tend to reside within the same loss basin, enabling linear interpolation between solutions without traversing high-loss regions. Singh and Jaggi (2020) demonstrated that neural networks sharing early training trajectories exhibit this connectivity property, while (Izmailov et al., 2018) provided theoretical analysis connecting proximity to initialization with favorable generalization behavior. Importantly, this connectivity assumption breaks down when models originate from independent random initializations or when fine-tuning procedures cause parameters to diverge substantially from the shared pretrained manifold, resulting in interpolation paths that traverse high-loss barriers (Singh et al., 2024).
The Model Soups framework, introduced by Wortsman et al. (2022a), represents a principled advancement in weight-space averaging by systematically studying the aggregation of multiple fine-tuned checkpoints. The framework establishes two primary aggregation strategies: the uniform soup and the greedy soup (Jang et al., 2024).
The uniform soup computes the arithmetic mean of all $N$ available model checkpoints without selection:
$\theta_{\text{uniform}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$
This baseline approach assumes all fine-tuned models contribute constructively to the aggregate solution. While computationally straightforward, the uniform soup may incorporate checkpoints whose parameter contributions introduce interference that degrades overall performance.
The greedy soup addresses this limitation through iterative conditional selection, incorporating models into the aggregate only when their inclusion improves held-out validation performance. Formally, starting with an initial model $\theta_{\text{soup}}=\theta_{1}$, each candidate $\theta_{i}$ is added only if:
$\mathcal{L}\left(\frac{k\cdot\theta_{\text{soup}}+\theta_{i}}{k+1}\right)<\mathcal{L}(\theta_{\text{soup}})$
where $k$ denotes the current number of models in the soup and $\mathcal{L}$ represents the validation loss function.
Empirically, Model Soups demonstrated substantial improvements across multiple benchmarks. While initially validated on vision models, this approach has since been applied to large language models. The key advantage is enhanced robustness to distribution shift. These improvements (Wortsman et al., 2022b) incur no additional inference cost compared to deploying a single model, as the merged architecture maintains identical computational requirements during deployment.
This selective aggregation strategy addresses a fundamental limitation of uniform averaging: not all models contribute equally to the merged solution, and certain checkpoints may introduce parameter interference that degrades performance. Wortsman et al. (2022a) demonstrate that weight-space operations can enable constructive task composition when models are appropriately combined, while unstructured aggregation frequently results in destructive interference patterns. Model Ratatouille (Ram’e et al., 2023a) extends this insight by demonstrating that recycling diverse auxiliary models fine-tuned on various domains improves out-of-distribution generalization, providing a principled approach for leveraging the growing ecosystem of available fine-tuned models. The relationship between model aggregation and federated learning has been explored extensively, with Li et al. (2023a) examining how weighted averaging strategies can be adapted to scenarios involving heterogeneous client contributions (Li et al., 2020; Wang et al., 2020) and non-identically distributed data.
Unlike exhaustive combinatorial search over model subsets, which becomes computationally prohibitive as the candidate pool grows exponentially, greedy selection provides an efficient approximation that scales linearly with the number of available checkpoints. However, this approach assumes that validation performance serves as a reliable proxy for generalization, an assumption that may not hold under significant distribution shift between validation and deployment domains. Furthermore, theoretical analyses indicate that linear averaging methods, while computationally efficient, do not account for the permutation symmetries inherent in neural network weight spaces, motivating the development of alignment-based interpolation techniques discussed in later subsections. Beyond these structural considerations, the uniform weighting scheme inherent in these approaches raises a more fundamental question: whether treating all parameters as equally important accurately reflects their heterogeneous contributions to model functionality.
3.2 Importance-Weighted Averaging and Fisher Information
Importance-weighted averaging methods directly address this question by rejecting the implicit assumption that all parameters contribute equally to model functionality, an assumption that significantly misrepresents the heterogeneous nature of neural network weight spaces. In practice, certain parameters encode critical task-specific information while others remain largely unchanged from pretraining or exhibit high variance across fine-tuning runs. By incorporating second-order information about parameter sensitivity, these methods enable more principled aggregation that prioritizes high-information parameters while attenuating noise from less consequential weights.
The theoretical foundation for importance-weighted merging derives from the Fisher Information Matrix (FIM), which characterizes the local curvature of the loss landscape regarding model parameters. For a model with parameters $\theta$ and loss function $\mathcal{L}$, the Fisher Information Matrix is defined as $F=\mathbb{E}[\nabla_{\theta}\log p(x|\theta)\nabla_{\theta}\log p(x|\theta)^{\top}]$, where the expectation is taken over the data distribution. Parameters associated with high Fisher information values correspond to directions in weight space where small perturbations significantly affect model predictions, indicating their importance for task performance. Fisher-weighted averaging exploits this insight by computing the merged parameters as $\theta_{\text{merged}}=(\sum_{i}F_{i})^{-1}\sum_{i}F_{i}\theta_{i}$, where $F_{i}$ denotes the Fisher Information Matrix computed for model $i$. This formulation effectively weights each model’s contribution inversely proportional to its uncertainty, yielding a maximum a posteriori estimate under Gaussian assumptions about parameter distributions.
The computational demands of full Fisher Information Matrix estimation render it intractable for contemporary large language models with billions of parameters. Practical implementations therefore employ diagonal approximations, where only the diagonal entries $F_{i,jj}$ are retained, reducing both memory requirements and computational complexity from $O(d^{2})$ to $O(d)$ where $d$ represents the parameter count. This diagonal Fisher-weighted averaging computes element-wise weighted means: $\theta_{\text{merged},j}=(\sum_{i}F_{i,jj}\theta_{i,j})/(\sum_{i}F_{i,jj})$, treating each parameter independently. While this approximation discards correlation structure between parameters, empirical investigations demonstrate that diagonal Fisher weighting substantially outperforms uniform averaging across diverse task combinations (Matena and Raffel, 2022).
Alternative formulations leverage covariance statistics rather than Fisher information to estimate parameter importance. The RegMean approach (Jin et al., 2023) computes inner products between model activations (Matena and Raffel, 2022) to construct weight matrices that account for the representational structure learned by each model. Specifically, RegMean solves a linear regression problem that minimizes the expected squared difference between the merged model’s outputs and those of the constituent models, weighted by activation covariances. This representation-aware weighting implicitly captures parameter dependencies that diagonal Fisher approximations neglect, potentially yielding superior merging outcomes when models have learned complementary feature representations. WiSE-FT (Wortsman et al., 2022b) demonstrates that robust fine-tuning of zero-shot models can be achieved through simple weight interpolation between the pretrained and fine-tuned parameters, providing a principled approach for preserving zero-shot capabilities while acquiring task-specific expertise. Bayesian perspectives on weighted aggregation naturally extend these importance-based schemes, treating merged parameters as posterior estimates given observations from multiple training procedures. Recent work has also explored the relationship between model merging and continual learning (He et al., 2024; Jordan et al., 2022), with Stoica et al. (2023) proposing ZipIt! for merging models trained on disjoint tasks by exploiting feature correlations, and Tang et al. (2024a) providing systematic benchmarking of merging methods under controlled conditions.
The relationship between model merging and catastrophic forgetting (Kirkpatrick et al., 2016) presents another critical consideration. When fine-tuned models are merged, the resulting model must retain capabilities from all constituent tasks without catastrophically forgetting any individual specialization. Elastic Weight Consolidation (EWC) and related continual learning methods provide theoretical frameworks for understanding which parameters are most critical for specific capabilities, insights that directly inform importance-weighted merging strategies.
Despite their theoretical appeal, importance-weighted methods introduce several practical challenges. Fisher information estimates require access to representative data samples from each constituent model’s training distribution, information that may be unavailable when merging pretrained models obtained from external sources. Furthermore, the approximation quality of diagonal Fisher estimates degrades in highly overparameterized networks where off-diagonal correlations capture essential structural relationships (Matena and Raffel, 2022). Unlike uniform averaging, which requires only the model weights themselves, importance-weighted approaches demand additional computational overhead for Fisher or covariance estimation, creating tradeoffs between merging quality and practical deployability that practitioners must navigate based on their specific resource constraints and accuracy requirements. Crucially, both uniform and importance-weighted averaging share a common assumption: they operate on models trained independently, treating the aggregation problem as combining discrete endpoints without exploiting any structure from the optimization process itself.
Limitations and failure modes.
Despite its simplicity and empirical success, linear averaging exhibits well-characterized failure modes. When source models are fine-tuned with substantially different learning rates or regularization strengths, the resulting parameter vectors may reside in distinct loss basins, causing their average to land on a high-loss barrier (Frankle et al., 2020). Performance also degrades when the number of merged models grows large and the tasks are semantically diverse, a phenomenon termed task interference. Ramé et al. (Ram’e et al., 2023a) partially address this by recycling intermediate fine-tuned checkpoints from diverse auxiliary tasks (Model Ratatouille), demonstrating that averaging models exposed to a broader training distribution yields more robust merged solutions. Nevertheless, characterizing the precise conditions under which linear averaging provably succeeds remains an open theoretical challenge (see Section 2).
3.3 Trajectory-Based Averaging and Optimization Dynamics
Trajectory-based averaging represents a significantly distinct paradigm that directly exploits the temporal structure of optimization, precisely the information that endpoint-focused methods discard. Rather than combining final checkpoints of separate training procedures, these techniques aggregate intermediate states collected along the optimization trajectory, leveraging the insight that weight-space averages of sequential iterates can approximate solutions residing near the center of flat loss basins. This geometric centering yields models exhibiting enhanced robustness to distribution shifts and improved generalization characteristics.
Polyak Averaging. The theoretical foundations for trajectory-based averaging trace to Polyak and Juditsky’s seminal work on iterate averaging for stochastic approximation. Their analysis established that averaging SGD iterates achieves optimal asymptotic convergence rates, with the averaged solution $\bar{\theta}_{n}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}$ converging at rate $O(1/n)$ compared to $O(1/\sqrt{n})$ for the final iterate alone. This classical result provides the theoretical underpinning for modern trajectory averaging methods, demonstrating that temporal averaging serves as an implicit regularizer that suppresses gradient noise while preserving convergence to stationary points.
Stochastic Weight Averaging. Stochastic Weight Averaging (SWA) adapts iterate averaging to modern deep learning by averaging checkpoints collected during the latter stages of training with cyclic or high constant learning rates (Izmailov et al., 2018). Formally, given checkpoints $\theta_{1},\theta_{2},\ldots,\theta_{n}$ collected after initial convergence, SWA computes $\theta_{\text{SWA}}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}$. The key insight is that SGD with appropriately tuned learning rates explores a connected region corresponding to a single basin of attraction, and averaging over this exploration identifies the basin’s centroid. In the context of LLMs, SWA has been successfully applied to smooth out the optimization trajectory during the final stages of pretraining or instruction tuning, leading to models with improved generation quality and reduced perplexity on held-out data. The flatness of these averaged solutions, measured via local entropy metrics, correlates strongly with improved generalization.
SWAG and Uncertainty Quantification. Stochastic Weight Averaging Gaussian (SWAG) extends SWA to approximate Bayesian inference by additionally capturing the covariance structure of the optimization trajectory. Beyond computing the mean $\theta_{\text{SWA}}$, SWAG maintains a low-rank approximation to the covariance matrix $\Sigma\approx\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{i}-\theta_{\text{SWA}})(\theta_{i}-\theta_{\text{SWA}})^{\top}$, enabling posterior sampling for uncertainty estimation. This extension proves particularly valuable for calibrated predictions and out-of-distribution detection, achieving competitive performance with ensemble methods at substantially reduced computational cost.
Lookahead Optimizer. The Lookahead optimizer provides a complementary perspective by integrating trajectory averaging directly into the optimization procedure rather than applying it post-hoc. Lookahead maintains slow weights $\phi$ updated every $k$ steps toward the fast weights $\theta$ via $\phi\leftarrow\phi+\alpha(\theta-\phi)$, effectively performing online trajectory averaging. This approach reduces variance in gradient estimates and demonstrates improved stability across learning rate choices, often achieving faster convergence while maintaining or improving final performance.
Exponential Moving Average. Exponential Moving Average (EMA) represents a continuously updated variant that maintains a shadow copy of parameters throughout training. At each iteration $t$, the EMA parameters are updated according to $\theta_{\text{EMA}}^{(t)}=\alpha\theta_{\text{EMA}}^{(t-1)}+(1-\alpha)\theta^{(t)}$, where $\alpha$ controls the temporal smoothing strength. Unlike SWA’s uniform averaging, EMA implements exponentially-decaying weights prioritizing recent iterates while retaining memory of the entire trajectory. The selection of decay coefficient $\alpha$ varies systematically with model scale and training duration. For large language models, values close to 1 are typical, with larger models and longer training runs generally benefiting from stronger smoothing.
Applications in LLM Training. Trajectory averaging has demonstrated substantial benefits in large-scale language model development. In neural machine translation and general LLM pretraining, checkpoint averaging of the final training checkpoints is commonly used (Izmailov et al., 2018). In the model merging context, averaging multiple checkpoints from the same fine-tuning run can improve generation quality and robustness on downstream NLP tasks (Wortsman et al., 2022a).
The connection between trajectory-based averaging and broader model merging frameworks emerges when considering multiple training runs initialized from common starting points. When models share a pretrained initialization and undergo fine-tuning on related tasks, their optimization trajectories remain confined to overlapping basins of attraction, enabling checkpoints to be meaningfully averaged despite originating from nominally independent procedures. This perspective bridges trajectory-based methods with task vector approaches, as temporal averaging within a single fine-tuning run and spatial averaging across related runs both operate by centering solutions within shared loss basins, with combined approaches yielding additive benefits.
Practical implementations of trajectory-based averaging require attention to checkpoint selection strategy and computational overhead. Regarding collection frequency, Izmailov et al. established that averaging every epoch during the final 25% of training provides near-optimal results, while more frequent collection yields diminishing returns (Izmailov et al., 2018). For memory efficiency, running average formulations avoid storing multiple checkpoints: SWA requires only two model copies (current and averaged), while EMA maintains constant memory overhead regardless of training duration. The robustness benefits to distribution shift have been validated in multiple studies, with SWA-averaged models exhibiting smaller performance degradation on domain-shifted evaluation sets compared to single checkpoints, attributed to the implicit regularization toward flatter minima.
Despite their effectiveness, trajectory-based methods share a common characteristic with the importance-weighted approaches discussed earlier: they perform arithmetic averaging in weight space, implicitly treating the parameter manifold as Euclidean. This assumption, while reasonable for iterates within a single optimization trajectory that remain geometrically proximate, becomes increasingly problematic when averaging models that have diverged substantially during independent fine-tuning, a limitation that motivates geometric interpolation techniques designed to preserve the statistical properties of constituent parameter sets.
Limitations.
Trajectory-based methods significantly require access to intermediate checkpoints from a single training run, making them inapplicable when only final model weights are available—the common scenario in open-weight model repositories. Furthermore, the memory overhead of storing $T$ checkpoints ($T\times|\theta|$ parameters) can be prohibitive for billion-parameter models, though rank-1 covariance approximations such as those in SWAG partially mitigate this.
3.4 Geometric Interpolation and Manifold-Aware Methods
Since Euclidean assumptions become problematic for divergent models, geometric interpolation techniques offer principled alternatives that account for the curved structure of parameter spaces. When combining two parameter vectors $\theta_{A}$ and $\theta_{B}$ via linear interpolation $\theta_{\text{merged}}=(1-\alpha)\theta_{A}+\alpha\theta_{B}$, the resulting vector’s magnitude typically satisfies $\|\theta_{\text{merged}}\|<(1-\alpha)\|\theta_{A}\|+\alpha\|\theta_{B}\|$ unless the vectors are perfectly aligned, leading to systematic magnitude shrinkage that can degrade model representations. This phenomenon proves particularly acute when merging models with substantially different weight configurations.
Spherical Linear Interpolation (SLERP) addresses magnitude preservation by constraining the interpolation path to lie on the hypersphere connecting two points in weight space. For parameter vectors normalized to unit length, SLERP computes the interpolated point as:
$\text{SLERP}(\theta_{A},\theta_{B};\alpha)=\frac{\sin((1-\alpha)\Omega)}{\sin(\Omega)}\theta_{A}+\frac{\sin(\alpha\Omega)}{\sin(\Omega)}\theta_{B}$
where $\Omega=\arccos(\theta_{A}\cdot\theta_{B}/\|\theta_{A}\|\|\theta_{B}\|)$ represents the angular distance between the two vectors. This formulation ensures that the interpolated vector maintains constant magnitude throughout the interpolation trajectory, traversing a geodesic arc rather than cutting through the interior of the sphere. Empirical investigations have demonstrated that SLERP-based merging preserves representational geometry more faithfully than linear averaging, particularly for attention weight matrices where directional information carries semantic significance.
In practice, applying SLERP to neural network merging requires careful consideration of granularity. Layer-wise application, where SLERP operates independently on each layer’s parameter tensor, has emerged as the dominant approach, as global application across flattened parameter vectors fails to capture the hierarchical structure of deep networks. observed that layer-wise (Yang et al., 2024c) SLERP achieves superior capability retention compared to uniform linear averaging when merging models fine-tuned on semantically distant tasks, attributing this advantage to the preservation of layer-specific activation statistics.
Beyond spherical interpolation, manifold-aware methods encompass techniques that explicitly model the non-Euclidean structure of parameter spaces. Latent space interpolation approaches, as explored in (Li et al., 2023a), project weight configurations into lower-dimensional manifolds where interpolation can be performed with reduced interference before mapping back to the original parameter space. This paradigm acknowledges that the effective dimensionality of viable weight configurations is substantially lower than the ambient parameter count, enabling interpolation paths that remain within the manifold of high-performing solutions.
The variance preservation problem presents additional challenges beyond magnitude shrinkage. When averaging independent parameter vectors, the variance of the resulting distribution decreases proportionally to the number of models merged, potentially collapsing learned features toward uninformative mean values. Rescaling strategies that adjust merged parameters to match the original variance statistics partially address this limitation, though they introduce hyperparameters governing the appropriate variance target. Recent work notes that geometric interpolation methods incorporating both magnitude and variance preservation consistently outperform naive averaging when source models exhibit low angular similarity, though the benefits diminish for models sharing high linear mode connectivity (Singh and Jaggi, 2020) where Euclidean assumptions remain approximately valid. While these geometric considerations inform how interpolation should proceed, an orthogonal question concerns what exactly should be interpolated, a perspective that motivates reconceptualizing fine-tuning through the lens of task vectors. Table 3 compares the weight-space averaging and geometric interpolation methods discussed above.
Table 3: Comparison of weight-space averaging and geometric interpolation methods for model merging.
Method Key Approach Key Feature(s) Limitations/Notes Year
Uniform Averaging
(Model Soups)
Element-wise arithmetic mean: $\theta_{\text{merged}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$ Simple; no hyperparameters; treats weight space as Euclidean Assumes equal parameter importance; suboptimal when models diverge significantly 2022
Fisher-Weighted Averaging Importance weighting via Fisher Information matrix Accounts for heterogeneous parameter importance; preserves critical task-specific weights Computationally expensive; requires data for Fisher estimation 2022
Stochastic Weight Averaging
(SWA) (Izmailov et al., 2018)
Average checkpoints along optimization trajectory Exploits temporal structure; finds flatter minima; improves generalization Limited to single training run; requires checkpoint storage 2018
SLERP Spherical linear interpolation on hypersphere Preserves weight vector magnitude; avoids norm shrinkage in linear interpolation Limited to pairwise merging; requires careful normalization 2023
Task Arithmetic
(Ilharco et al., 2023)
Compute and combine task vectors $\tau=\theta_{\text{ft}}-\theta_{\text{pre}}$ Modular; supports task addition, negation, and composition Task interference; requires shared pretrained initialization 2023
TIES-Merging
(Yadav et al., 2023)
Trim redundant params, elect signs, merge disjoint sets Reduces parameter interference; resolves sign conflicts Hyperparameter sensitive (trim ratio); additional computation 2023
Latent Space Merging
(Li et al., 2023a)
Interpolation in activation/latent space rather than weights Bypasses weight space geometry issues; often outperforms weight averaging May require data access; increased inference complexity 2024
Manifold-Aware Interpolation Geodesic paths on parameter manifolds with permutation alignment Respects non-Euclidean geometry; handles mode connectivity Computationally intensive; requires solving alignment problem 2023
The methods examined in this section demonstrate that weight-space averaging, whether through uniform linear combinations, Fisher-weighted importance schemes, trajectory-based accumulation along optimization paths, or geometry-aware interpolation on parameter manifolds, provides a powerful framework for combining multiple models without additional training, with performance gains stemming from the exploitation of loss landscape connectivity and mode connectivity properties. While these approaches focus on holistically merging entire model checkpoints through various averaging strategies, the next section examines task vector arithmetic and sparsification-enhanced methods, which shift the perspective from combining full parameter sets to treating fine-tuning-induced weight differences as modular, composable units supporting addition, negation, and scaling for targeted multi-task composition.
4 Task Vector Arithmetic and Sparsification-Enhanced Methods
We now turn to a different paradigm: isolating and manipulating the task-specific knowledge encoded during fine-tuning. Task vector arithmetic, introduced by Ilharco et al. (2023), reconceptualizes model merging by treating the weight differences between fine-tuned models and their shared pre-trained ancestor as composable vectors that support algebraic operations such as addition, negation, and scaling. We also cover sparsification-enhanced variants such as TIES-Merging (Yadav et al., 2023) and DARE (Yu et al., 2023) that address parameter interference and sign conflicts.
4.1 Task Vector Formulation and Conceptual Framework
The task vector framework formalizes how fine-tuning encodes task-specific knowledge within neural network parameters. Ilharco et al. (2023) introduced this foundational framework by proposing that the knowledge acquired during fine-tuning can be isolated, manipulated, and transferred through simple algebraic operations on model weights. This formulation enables model specialization becomes amenable to systematic composition and modification.
Formally, given a pretrained model with parameters $\theta_{\text{pre}}$ and a model fine-tuned on task $t$ with resulting parameters $\theta_{t}$, the task vector $\tau_{t}$ is defined as the element-wise difference:
$\tau_{t}=\theta_{t}-\theta_{\text{pre}}$ (3)
This definition captures the parametric displacement induced by fine-tuning, effectively encoding the directional shift in weight space that transforms a general-purpose foundation model into a task-specialized variant. Based on this formulation, Ilharco et al. (2023) defined three fundamental arithmetic operations that constitute a versatile toolkit for training-free model editing:
Task Addition (Multi-task Fusion). To combine capabilities from multiple tasks, their respective task vectors can be added and applied to the pretrained model:
$\theta_{\text{multi}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$ (4)
where $\lambda_{i}$ denotes the scaling coefficient for task $i$. This operation enables the creation of multi-task models without joint training, assuming the task vectors reside in approximately orthogonal subspaces.
Task Negation (Forgetting/Unlearning). To remove an undesirable capability or behavior (e.g., toxicity) encoded in a task vector $\tau$, the vector can be subtracted from the pretrained model:
$\theta_{\text{forget}}=\theta_{\text{pre}}-\lambda\tau$ (5)
where $\lambda>0$ controls the strength of the unlearning effect. This operation has found practical application in detoxification and bias removal for LLMs.
Task Scaling (Capability Modulation). The strength of a specific capability can be continuously modulated by scaling its task vector:
$\theta_{\text{scaled}}=\theta_{\text{pre}}+\alpha\tau$ (6)
where $\alpha$ controls the intensity of the specialized behavior ($\alpha>1$ amplifies the capability, while $0<\alpha<1$ attenuates it).
Task Analogy (Relationship Transfer). Extending the vector algebra metaphor, task analogy enables transferring the relationship between two tasks to a third domain. Given task vectors $\tau_{A}$, $\tau_{B}$, and a target task vector $\tau_{C}$, the analogy operation applies the directional difference between two tasks to a third:
$\theta_{\text{analogy}}=\theta_{\text{pre}}+\tau_{C}+(\tau_{B}-\tau_{A})$ (7)
This operation mirrors classical word analogy reasoning (e.g., “king - man ++ woman == queen”) in embedding spaces, but operates directly on model parameters. As illustrated in Figure 3(d), task analogy enables compositional transfer of inter-task relationships without additional training.
The conceptual elegance of this formulation lies in its interpretation: fine-tuning is reconceptualized not as an opaque optimization process but as the application of a structured transformation that can be decomposed, analyzed, and recombined with other such transformations.
The task vector paradigm derives its power from treating these parameter deltas as first-class mathematical objects inhabiting a vector space (Ilharco et al., 2023). Under this interpretation, the weight space of neural networks admits linear algebraic operations that preserve semantic meaning. When a model is fine-tuned on a specific capability, whether instruction following, mathematical reasoning, or domain expertise, the resulting task vector encodes a compact representation of that capability that remains meaningful when extracted from its original context and applied to compatible base models.
This formulation maintains deep connections to the theoretical foundations of mode connectivity discussed in earlier sections. The success of task vector arithmetic implicitly relies on the phenomenon of linear mode connectivity (Singh and Jaggi, 2020) between fine-tuned models sharing a common pretrained initialization (Garipov et al., 2018). When fine-tuned models reside within the same loss basin as their shared pretrained ancestor, the linear interpolation paths between them traverse regions of relatively low loss, enabling meaningful combination through arithmetic operations on their corresponding task vectors.
The requirements for task vector compatibility impose important constraints on practical applications. As emphasized in (Li et al., 2023a), task vector operations are inherently limited to models with identical architectures since they rely on element-wise parameter correspondence. Both source models must derive from the same pretrained checkpoint, ensuring that their parameter spaces are aligned and that the extracted task vectors encode semantically coherent transformations. This architectural constraint reflects the fundamental observation that task vectors represent relative displacements within a shared reference frame established by the common initialization.
The magnitude and distribution of task vector components carry meaningful information about the nature of fine-tuning adaptations. Ilharco et al. (2023) investigated how different fine-tuning procedures produce task vectors with distinct statistical properties, affecting downstream merging behavior. Empirical analyses reveal that task vectors exhibit considerable sparsity, with relatively few parameters undergoing substantial modification during fine-tuning while the majority remain close to their pretrained values (Yu et al., 2023). This sparsity observation has motivated various enhancement strategies, though the immediate practical power of task vectors emerges from the fundamental algebraic operations they support, addition, negation, and scaling, which together constitute a versatile toolkit for training-free model editing. Similar to how knowledge distillation (Kim et al., 2023; Hinton et al., 2015) transfers knowledge between models through output matching, task vectors enable knowledge transfer through direct parameter manipulation. Figure 3 illustrates the three primary task vector operations.
(a) Addition
$\tau_{A}$$\tau_{B}$$\theta_{\text{merged}}$
$\theta_{\text{pre}}+\lambda(\tau_{A}+\tau_{B})$
(b) Negation
$\theta_{\text{pre}}$$\tau$$\theta_{\text{clean}}$
$\theta_{\text{pre}}-\lambda\tau_{\text{toxic}}$
(c) Scaling
$\theta_{\text{pre}}$$\lambda\tau$$\theta_{\text{merged}}$++
$\theta_{\text{pre}}+\lambda\tau$, $\lambda\in[0,1]$
(d) Analogy
$\tau_{A}$$\tau_{B}$$\theta^{\prime}_{\text{new}}$-apply
$\theta_{\text{new}}+(\tau_{B}-\tau_{A})$
Figure 3: Four fundamental task vector operations. (a) Addition combines multiple task vectors into a single model. (b) Negation subtracts a task vector to remove unwanted behavior. (c) Scaling controls task influence via $\lambda$. (d) Analogy transfers the relationship between two tasks.
4.2 Basic Arithmetic Operations for Model Editing
The three primary operations—addition, negation, and scaling—transform task vectors into versatile tools for training-free model editing.
Task Vector Addition represents the most widely employed operation, enabling the combination of multiple specialized capabilities into a unified model. Given a set of $n$ task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ extracted from models fine-tuned on distinct tasks, the merged model parameters are computed as $\theta_{\text{merged}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$, where $\lambda_{i}$ denotes the scaling coefficient for task $i$. The seminal work Ilharco et al. (2023) demonstrated that this simple additive combination successfully transfers capabilities across diverse task categories, including instruction following, mathematical reasoning, and domain-specific knowledge (Xiao et al., 2023) adaptation in LLMs. The effectiveness of addition relies critically on the assumption that task-specific knowledge encoded in different task vectors occupies approximately orthogonal subspaces of the parameter manifold, allowing superposition without destructive interference. (Li et al., 2023a) provides comprehensive empirical evidence that task addition achieves competitive performance across numerous NLP benchmarks when source tasks exhibit complementary rather than conflicting requirements.
Task Vector Negation offers an effective capability for targeted behavior removal without requiring access to negative training examples or explicit unlearning procedures. By computing $\theta_{\text{edited}}=\theta_{\text{pre}}-\lambda\tau_{\text{unwanted}}$, practitioners can attenuate or eliminate specific model behaviors such as toxic language generation, demographic biases, or domain-specific knowledge (Xiao et al., 2023) that should be restricted. Ilharco et al. (2023) originally demonstrated this capability by showing that negating task vectors extracted from models fine-tuned on toxic content datasets reduced harmful generation rates while preserving general linguistic competence. The mathematical intuition underlying negation follows from the interpretation of task vectors as directional displacements in capability space; reversing the direction of displacement moves the model away from the associated capability manifold. However, Ilharco et al. (2023) revealed important nuances regarding fairness implications, demonstrating that naive negation operations can inadvertently introduce new biases or disproportionately affect model behavior across different demographic groups. This finding underscores the need for careful analysis when employing negation for bias mitigation purposes.
Scaling Coefficients provide fine-grained control over the magnitude of task contributions, enabling balance competing objectives and prevent any single task from dominating the merged representation. The choice of scaling parameter $\lambda$ presents a fundamental tradeoff: larger values increase task-specific performance but risk catastrophic interference with the pretrained model’s general capabilities, while smaller values preserve base model competence at the cost of diminished task specialization. Unlike Ilharco et al. (2023), which primarily explored uniform scaling across tasks, subsequent investigations have demonstrated the benefits of task-specific coefficient tuning. Tang et al. (2024c) showed that operating in the tangent space of neural network weights enables more principled coefficient selection by linearizing the relationship between scaling magnitude and performance outcomes.
The analogy between task arithmetic and classical linear algebra operations provides both conceptual clarity and practical limitations. While vector addition in Euclidean space enjoys exact superposition properties, task vector addition operates in a highly nonlinear parameter space where interactions between components can produce emergent behaviors not predicted by simple linear models  (Li et al., 2023a). These nonlinear interactions manifest as parameter-level conflicts that can substantially degrade merged model performance, interference phenomena that demand systematic examination to enable principled mitigation strategies.
4.3 Parameter Interference and Sign Conflict Analysis
The parameter-level conflicts identified in the preceding discussion manifest through three distinct interference mechanisms that emerge when multiple task vectors are combined through element-wise operations. Systematic examination of these mechanisms is essential for developing principled mitigation strategies that preserve the compositional benefits of task arithmetic while minimizing performance degradation.
Sign Conflicts represent perhaps the most pernicious form of parameter interference, occurring when task vectors from different fine-tuning objectives contain opposing directional updates for identical weight positions. Formally, for task vectors $\tau_{1},\tau_{2},\ldots,\tau_{n}$ derived from $n$ distinct fine-tuning procedures, a sign conflict at parameter position $j$ occurs when $\exists i,k:\text{sign}(\tau_{i}^{(j)})\neq\text{sign}(\tau_{k}^{(j)})$ and both $|\tau_{i}^{(j)}|,|\tau_{k}^{(j)}|>\epsilon$ for some threshold $\epsilon$. When such conflicts arise, naive averaging produces parameter values that represent a compromised intermediate state, potentially satisfying neither task’s requirements. Yadav et al. (2023) provided comprehensive empirical analysis demonstrating that sign conflicts correlate strongly with performance degradation in merged models, with certain layer types exhibiting systematically higher conflict rates than others. The analysis suggests that attention projection matrices and feed-forward intermediate layers may accumulate more severe conflicts, indicating that these components encode task-specific information that resists direct composition.
Magnitude Disparities constitute a second critical interference mechanism, arising from the heterogeneous scaling of parameter updates across different fine-tuning procedures. Even when sign agreement exists, substantial differences in update magnitudes, quantified as $\max_{i}|\tau_{i}^{(j)}|/\min_{k}|\tau_{k}^{(j)}|$ for non-trivial updates, can cause certain tasks to dominate the merged representation disproportionately. This phenomenon is particularly pronounced when combining models fine-tuned with different learning rates, dataset sizes, or training durations (Yadav et al., 2023). The resulting merged model exhibits biased behavior favoring tasks whose fine-tuning produced larger magnitude updates, effectively drowning out the contributions of more conservatively trained variants. Subsequent investigations by Li et al. (2023a) demonstrate that real-world task vectors exhibit heavy-tailed magnitude distributions with significant inter-task variance.
Redundant Parameter Modifications introduce a subtler but equally important source of interference, wherein low-magnitude updates that do not encode meaningful task-specific knowledge accumulate noise during the merging process. These redundant parameters arise from optimization dynamics that produce non-zero gradients even for weights irrelevant to the target task, creating spurious modifications that dilute the signal-to-noise ratio of the merged representation. Yu et al. (2023) formalized this observation by analyzing the effective rank and information content of task vectors, finding that a substantial fraction of parameter updates contribute minimally to task performance while actively degrading merge quality through interference with truly informative modifications.
The interaction between these interference mechanisms produces complex, often non-additive effects that resist simple analytical characterization. (Li et al., 2023a) introduced statistical frameworks for analyzing the joint distribution of task vector parameters, revealing that interference severity depends critically on the correlation structure between task vectors rather than their marginal statistics alone. Visualization techniques including parameter-wise conflict heatmaps and layer-wise interference scores have emerged as diagnostic tools for identifying problematic regions before merging. This comprehensive characterization of interference phenomena, spanning sign conflicts, magnitude disparities, and redundant modifications, directly motivates the sparsification-enhanced approaches that selectively retain informative parameters while eliminating conflict-inducing modifications.
4.4 Sparsification-Enhanced Merging Methods
Sparsification-enhanced methods address these interference phenomena by exploiting redundancy in task vectors: strategic parameter elimination reduces conflicts while preserving task-critical information. These approaches transform dense task vectors into sparse representations with improved composability.
The seminal Trim-Elect-Sign (TIES) mechanism introduced a three-stage pipeline that systematically addresses each category of interference. Given task vectors $\{\tau_{i}\}_{i=1}^{n}$ derived from $n$ fine-tuned models, the trimming stage first eliminates parameters with magnitudes below a threshold $\lambda$, producing sparse vectors $\tilde{\tau}_{i}$ where $\tilde{\tau}_{i,j}=\tau_{i,j}\cdot\mathbf{1}$. The election stage then resolves sign conflicts through majority voting, computing a consensus sign mask $\gamma_{j}=\text{sgn}(\sum_{i}\tilde{\tau}_{i,j})$ that determines the dominant direction for each parameter across all task vectors. Finally, the sign alignment stage retains only those parameter updates consistent with the elected sign, yielding conflict-free task vectors suitable for aggregation.
Algorithm 1 TIES-Merging: Trim, Elect Sign, and Merge
0: Task vectors $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$, trim ratio $k$, pretrained weights $\theta_{\text{pre}}$
0: Merged model parameters $\theta_{\text{merged}}$
1:// Step 1: Trim - Keep only top-kk% parameters by magnitude
2:for $i=1$ to $n$ do
3:  $\tilde{\tau}_{i}\leftarrow\text{TopK}(\tau_{i},k)$ {Zero out small-magnitude parameters}
4:end for
5:// Step 2: Elect Sign - Resolve sign conflicts via majority voting
6:for each parameter position $j$ do
7:  $\gamma_{j}\leftarrow\text{sign}\left(\sum_{i=1}^{n}\tilde{\tau}_{i,j}\right)$ {Determine dominant sign}
8:end for
9:// Step 3: Merge - Average parameters with elected signs
10:for each parameter position $j$ do
11:  $S_{j}\leftarrow\{i:\text{sign}(\tilde{\tau}_{i,j})=\gamma_{j}\}$ {Select agreeing task vectors}
12:  $\bar{\tau}_{j}\leftarrow\frac{1}{|S_{j}|}\sum_{i\in S_{j}}\tilde{\tau}_{i,j}$ {Average aligned values}
13:end for
[…]
15:return $\theta_{\text{merged}}$
Yadav et al. (2023) demonstrated that this mechanism achieves substantial performance improvements over naive task arithmetic, particularly when merging more than three task vectors where sign conflicts proliferate.
Complementing the deterministic TIES approach, the DARE (Drop And REscale) framework (Yu et al., 2023) introduced probabilistic sparsification as an alternative conflict resolution strategy. Rather than employing magnitude-based pruning, DARE applies random dropout to task vector parameters with probability $p$, followed by rescaling of the retained parameters by factor 1/(1p)1/(1-p) to preserve the expected contribution. Mathematically, the sparsified task vector becomes $\hat{\tau}_{i}=\frac{1}{1-p}\cdot\tau_{i}\odot m_{i}$ where $m_{i}$ is a binary mask with $m_{i,j}\sim\text{Bernoulli}(1-p)$. This approach draws theoretical justification from the observation that fine-tuning induces primarily low-rank modifications to pretrained weights, similar to findings in network pruning research (Frankle and Carbin, 2019), implying substantial redundancy that can be exploited through random sampling without catastrophic information loss. Recent extensions such as DELLA-Merging (Deep et al., 2024) further refine this approach through magnitude-based sampling (Deep et al., 2024) strategies. The stochastic nature of DARE provides an implicit ensemble effect when multiple merging trials are conducted, though practical deployments typically employ a single sparsification pass.
Building upon these foundational methods, importance-aware sparsification has been proposed that leverages second-order information to guide parameter retention decisions. Rather than treating all parameters equivalently during the trimming process, such methods compute importance scores based on the Fisher information or gradient magnitude accumulated during fine-tuning, preferentially retaining parameters whose modification contributed most significantly to task adaptation. This approach addresses a key limitation of magnitude-based pruning, wherein parameters with small absolute values may nonetheless encode critical task-specific information.
Subsequent work extended the sparsification paradigm by introducing conflict-awareness into the retention decision process. Unlike TIES, which considers sign conflicts only after independent trimming of each task vector, conflict-aware methods jointly optimize sparsification masks across all source models to minimize expected interference while maximizing information retention. These methods employ a balanced sparsification criterion that normalizes contribution magnitudes across tasks, preventing dominant task vectors from overwhelming weaker contributors during the merge.
Further refinements have incorporated distributional statistics of task vector parameters into sparsification decisions. Recognizing that optimal trimming thresholds vary across layers and parameter groups, these methods adaptively calibrate sparsification intensity based on local magnitude distributions, achieving more precise conflict resolution than global threshold approaches. While these sparsification techniques have proven effective at mitigating interference, they nonetheless operate within the constraints of direct weight-space manipulation, motivating alternative formulations that reconsider the mathematical foundations of task vector composition itself.
4.5 Advanced Task Vector Manipulation and Recent Extensions
Motivated by these limitations of direct weight-space manipulation, recent methodological advances have extended the task arithmetic framework toward more principled mathematical reformulations that address fundamental limitations of standard parameter arithmetic. A particularly influential direction involves reconceptualizing task vector operations within the tangent space of neural network parameters, motivated by the observation that standard weight-space arithmetic implicitly assumes local linearity that may not hold for large parameter displacements.
Research in this area introduced a linearized formulation that operates on task vectors projected into the tangent space centered at the pretrained model weights (Tang et al., 2024c). Given a pretrained model with parameters $\theta_{0}$ and a fine-tuned model $\theta_{t}$, the standard task vector $\tau_{t}=\theta_{t}-\theta_{0}$ captures parameter displacements without accounting for the nonlinear relationship between weights and model behavior. The tangent space approach instead considers the linearized model $f_{\text{lin}}(x;\theta)=f(x;\theta_{0})+\nabla_{\theta}f(x;\theta_{0})^{\top}(\theta-\theta_{0})$, which provides a first-order approximation of how parameter changes affect model outputs. Within this framework, task vector arithmetic operates on the linearized representations, yielding improved compositional properties when combining multiple task-specific adaptations. Empirically, tangent space task arithmetic demonstrates superior performance on model editing benchmarks, particularly for negation operations where the standard formulation often produces unstable results.
The theoretical motivation underlying tangent space methods connects to the broader question of why task arithmetic succeeds despite the highly nonlinear nature of neural network loss landscapes. The linearization effectively constrains operations to a regime where the superposition principle approximately holds, addressing scenarios where direct weight interpolation violates the implicit linearity assumptions of standard task arithmetic. This reformulation proves especially valuable for combining task vectors with large magnitudes or when source models have undergone extensive fine-tuning that displaces parameters significantly from the pretrained initialization.
Complementary to these mathematical reformulations, adaptive rank-based approaches have emerged as powerful techniques for optimizing the representational structure of task vectors. Recent work proposed methods that adaptively determine optimal rank configurations for low-rank approximations of task vectors across different layers and modules. Unlike uniform rank reduction strategies, these approaches recognize that different network components encode task-specific information with varying degrees of redundancy, allocating representational capacity accordingly. The method formulates rank selection as an optimization problem that balances compression efficiency against information preservation, leveraging singular value decomposition to identify the effective dimensionality of task-specific modifications.
The integration of adaptive rank methods with existing sparsification techniques yields synergistic benefits. While sparsification operates by zeroing individual parameters, rank-based approaches restructure the entire task vector representation, capturing different aspects of redundancy in fine-tuned models. Table 4 summarizes these methods. This complementarity enables hybrid strategies that first reduce task vector rank to eliminate correlated modifications, then apply sparsification to address residual conflicts among the compressed representations. Li et al. (2023a) provided comprehensive empirical analysis demonstrating that such combined approaches consistently outperform either technique in isolation across diverse task combinations.
From a practical implementation perspective, these advanced manipulation methods have been progressively incorporated into community toolkits that facilitate reproducible model merging experimentation. The modular architecture of modern merging frameworks enables researchers to compose tangent space projections, adaptive rank reduction, and conflict-aware sparsification into customized merging pipelines tailored to specific application requirements. This ecosystem integration significantly lowers barriers to adoption, allowing practitioners to leverage sophisticated mathematical formulations without implementing complex algorithms from scratch. Despite their increasing sophistication, however, all task vector manipulation techniques discussed thus far share a common assumption: that merging should ultimately produce a unified parameter set. An alternative paradigm instead preserves distinct expert components while enabling flexible, input-dependent utilization of specialized capabilities.
Table 4: Comparison of task vector arithmetic and sparsification-enhanced methods for model merging.
Method Key Approach Key Feature(s) Limitations/Notes Year
Task Arithmetic
(Ilharco et al., 2023)
Linear task vector combination Defines task vectors as weight differences; enables addition, negation, scaling operations Suffers from parameter interference and sign conflicts 2023
TIES-Merging
(Yadav et al., 2023)
Trim, Elect Sign, Merge Resolves sign conflicts via magnitude-based trimming and democratic sign election Requires hyperparameter tuning for sparsity threshold 2023
DARE Drop and Rescale Randomly drops delta parameters and rescales remaining ones to preserve expectations Stochastic nature may introduce variance 2024
CABS Conflict-Aware Balanced Sparsification Explicitly identifies and eliminates conflicting parameters while maintaining balance Limited to identical architectures 2024
Tangent Space Methods Tangent Space Operations Reformulates task arithmetic in tangent space for improved linearity assumptions Increased mathematical complexity 2024
Task vector arithmetic provides a powerful conceptual framework for model editing, enabling intuitive operations such as knowledge addition and negation, while sparsification-enhanced methods like TIES-Merging and DARE address critical challenges of parameter interference and sign conflicts by selectively pruning redundant or conflicting parameters before aggregation. These approaches operate primarily at the individual parameter level, treating model weights as independent quantities to be manipulated through algebraic operations. While this section focused on parameter-space manipulations and sparsity-driven conflict resolution, the next section examines structured and information-guided merging approaches that move beyond element-wise operations to exploit architectural organization through Mixture-of-Experts (Shazeer et al., 2017) compositions, leverage activation statistics for principled merging decisions, and employ evolutionary search to automatically discover optimal merging configurations across the full design space.
5 Structured and Information-Guided Merging Approaches
While task vector arithmetic and sparsification techniques offer powerful tools for mitigating parameter interference, they primarily operate on static weight representations without considering the underlying model architecture or runtime behavior. We now turn to a complementary class of approaches that leverage structural priors and dynamic information to guide the merging process more effectively. This section examines three advanced paradigms: architecture-level composition methods that preserve modularity through Mixture-of-Experts style routing mechanisms, activation-informed techniques that exploit representation statistics for principled parameter alignment, and evolutionary or search-based meta-strategies that automatically discover optimal merging configurations across the design space.
Table 5: Unified comparison of representative model merging methods. Data-Free: no training/calibration data required. Interference: explicit handling of parameter conflicts. Shared Init: requires common pretrained initialization.
Method Category Data-Free Interference Handling Shared Init Scalability Year
Weight-Space Averaging (Section 3)
Simple Averaging Static High
Model Soups
(Wortsman et al., 2022a)
Greedy selection High 2022
Fisher Merging
(Matena and Raffel, 2022)
Importance-weighted Partial Medium 2022
RegMean
(Jin et al., 2023)
Linear regression Partial Medium 2023
SLERP Geometric High 2023
Task Vector & Sparsification (Section 4)
Task Arithmetic
(Ilharco et al., 2023)
Additive High 2023
TIES-Merging
(Yadav et al., 2023)
Sign resolution High 2023
DARE
(Yu et al., 2023)
Sparsification High 2023
Model Breadcrumbs
(Yadav et al., 2024b)
Sparse masking High 2023
DELLA
(Deep et al., 2024)
Magnitude sampling High 2024
Model Stock
(Deep et al., 2024)
Geometric selection Implicit High 2024
AdaMerging
(Yang et al., 2024d)
Adaptive coefficients Medium 2024
Structured & Routing (Section 5)
Git Re-Basin
(Ainsworth et al., 2023)
Permutation alignment N/A Low 2023
ZipIt!
(Stoica et al., 2023)
Feature matching N/A Low 2023
Branch-Train-Merge
(Li et al., 2022)
MoE routing N/A High 2022
PHATGOOSE
(Muqeeth et al., 2024)
Learned routing N/A Medium 2024
Evol. Merging
(Akiba et al., 2024)
Search-based Implicit Medium 2024
LoRA (Hu et al., 2022)-Specific
LoRAHub
(Huang et al., 2023a)
Bayesian composition Partial High 2023
MoLE
(Wu et al., 2024)
Expert routing N/A High 2025
5.1 Mixture-of-Experts Style Merging and Expert Routing Strategies
Mixture-of-Experts (MoE) style merging preserves specialized capabilities by organizing distinct expert pathways, directly addressing destructive interference when combining divergent models (Li et al., 2022; Muqeeth et al., 2023). By maintaining separate parameter spaces for each expert and employing learned routing mechanisms, MoE-style merging achieves a principled balance between capability preservation and unified deployment.
The Branch-Train-Merge (BTM) framework provides the conceptual foundation for scalable expert training within this paradigm (Li et al., 2022). In BTM, a shared pretrained LLM is first replicated into multiple branches, each subsequently fine-tuned independently on distinct data subsets or task distributions. The resulting specialized models are then integrated through routing mechanisms that direct inputs to appropriate experts based on learned criteria. Formally, given a set of $K$ expert models $\{E_{1},E_{2},\ldots,E_{K}\}$ derived from a common pretrained initialization, the merged output for an input $x$ is computed as:
$y=\sum_{i=1}^{K}g_{i}(x)\cdot E_{i}(x)$ (8)
where $g_{i}(x)$ represents the routing weight assigned to expert $i$ by a gating network $G$, typically computed via a softmax transformation over learned routing logits:
$g_{i}(x)=\frac{\exp(w_{i}^{\top}h(x))}{\sum_{j=1}^{K}\exp(w_{j}^{\top}h(x))}$ (9)
with $h(x)$ denoting an intermediate hidden representation and $\{w_{1},\ldots,w_{K}\}$ the learnable routing parameters. To promote computational efficiency (Coste et al., 2024), top-$k$ sparse routing activates only a subset of experts per input:
$g_{i}^{\text{top-}k}(x)=\begin{cases}g_{i}(x)&\text{if }i\in\text{TopK}(\{g_{j}(x)\}_{j=1}^{K})\\ 0&\text{otherwise}\end{cases}$ (10)
enabling inference costs that scale as $O(k/K)$ relative to the full ensemble. CALM (Composition to Augment Language Models) (Bansal et al., 2024) shows that LLMs can be augmented through composition with specialized modules, enabling capability expansion without full model retraining. The WEMoE framework (Tang et al., 2024d) demonstrates how this weight-ensembling approach can effectively integrate multiple task-specific models into a unified MoE architecture, achieving superior multi-task performance compared to naive averaging strategies by constructing experts from merged weight differences rather than full model copies.
Several prominent methods have emerged for constructing MoE systems from independently trained models. PHATGOOSE (Post-Hoc Adaptive Tokenwise Gating Over an Ocean of Specialized Exp.[truncated] addresses the challenge of routing over pre-existing fine-tuned models by learning lightweight gating parameters while keeping expert weights frozen. For each expert $E_{i}$, PHATGOOSE computes routing scores based on the dot product between a learned gate vector $v_{i}$ and the input representation: $g_{i}(x)=v_{i}^{\top}h(x)$, where gate vectors are optimized on each expert’s original training data to identify inputs the expert handles well. MoLE (Mixture of LoRA Experts) (Hu et al., 2022; Feng et al., 2024) takes a parameter-efficient approach by treating individual LoRA adapters as experts, composing the output as $y=W_{0}x+\sum_{i=1}^{K}g_{i}(x)\cdot B_{i}A_{i}x$, where $W_{0}$ represents the frozen pretrained weights and $B_{i}A_{i}$ denotes the low-rank decomposition of expert $i$. This LoRA-as-experts paradigm substantially reduces memory overhead since each expert contributes only the low-rank adapter parameters rather than full model weights. LoRAHub (Huang et al., 2023a) enables efficient cross-task generalization via dynamic LoRA composition, while LoRAMoE (Dou et al., 2023) extends this concept by using LoRA-based experts as plugins that preserve world knowledge during instru.[truncated]
The design of routing mechanisms constitutes a critical component (Muqeeth et al., 2024) determining MoE effectiveness. Gating networks typically employ learned transformations that map input representations to probability distributions over experts. The softmax-based router represents the most common instantiation, computing routing weights as $g(x)=\text{softmax}(W_{g}\cdot h(x))$, where $h(x)$ denotes an intermediate representation and $W_{g}$ represents learnable gating parameters. To promote computational efficiency through sparse activation, top-$k$ routing selects only a subset of experts for each input: $g_{i}(x)=\text{softmax}(\text{TopK}(W_{g}\cdot h(x)))_{i}$, enabling inference costs that scale sublinearly with expert count. Load balancing mechanisms prevent expert collapse, where routing degenerates to selecting only a few experts. The auxiliary load balancing loss $\mathcal{L}_{\text{balance}}=\alpha\cdot K\cdot\sum_{i=1}^{K}f_{i}\cdot p_{i}$ penalizes uneven expert utilization, where $f_{i}$ represents the fraction of tokens routed to expert $i$ and $p_{i}$ denotes the average routing probability for expert $i$ across the batch (Jiang et al., 2024; Fedus et al., 2022). Switch Transformers (Fedus et al., 2022) demonstrated scaling to trillion parameter models with simple and efficient sparsity mechanisms. Branch-Train-MiX (Sukhbaatar et al., 2024) provides a framework for mixing expert LLMs into a unified MoE architecture through branched training procedures. Alternative routing strategies beyond learned softmax gating include expert-choice routing (Jiang et al., 2024), which inverts the selection process by having experts select their top-$k$ preferred tokens rather than tokens selecting experts, naturally achieving load balance. The granularity of routing decisions, whether at token-level for fine-grained adaptation or sequence-level for reduced overhead, depends substantially on expert specialization patterns and input diversity (Yadav et al., 2024a).
Post-hoc integration of specialized fine-tuned models as experts presents unique challenges distinct from training MoE systems from scratch, as experts were never optimized for selective activation. PHATGOOSE addresses this through a two-stage process: first computing per-expert gate vectors by optimizing routing scores on each expert’s training data independently, then combining these gates for joint inference without requiring access to all training data simultaneously. Model Kinship (Hu et al., 2024) provides a framework for exploring and quantifying the relationships between models that influence merging outcomes. Arrow (Yadav et al., 2024a) formulates expert selection as multi-objective optimization, learning to compose subsets of LoRA experts based on task characteristics. ZipIt (Stoica et al., 2023) demonstrates that models trained on different tasks can be merged without additional training through feature matching (Stoica et al., 2023) and permutation alignment. For scenarios where expert training data is unavailable, representation-based heuristics estimate input-expert compatibility through embedding similarity (Hu et al., 2024). Model-Based Experts (Yadav et al., 2024a) further improve efficiency by training small auxiliary classifiers to predict which experts will produce outputs most similar to the full ensemble, enabling dynamic expert pruning at inference time.
MoE-style merging incurs distinct resource tradeoffs compared to parameter-averaging approaches. While parameter averaging produces a single model with memory requirements identical to any constituent, MoE architectures must store all $K$ expert parameter sets, resulting in memory scaling of $O(K\cdot|E|)$ where $|E|$ denotes expert size, though LoRA-based experts substantially mitigate this overhead (Hu et al., 2022). Computationally, sparse top-$k$ routing reduces inference FLOPs to approximately $k/K$ of the dense ensemble cost, but introduces routing computation overhead and memory bandwidth costs from loading selected expert weights. In contrast, merged models via techniques such as TIES-Merging (Yadav et al., 2023) or DARE (Yu et al., 2023) incur no additional inference cost beyond a single model but cannot dynamically adapt computation to input characteristics. The choice between paradigms thus depends on deployment constraints: MoE architectures favor scenarios requiring preserved specialist capabilities and sufficient memory, while parameter averaging suits resource-constrained settings where moderate capability blending suffices. Table 6 compares representative MoE-style and LoRA-based composition methods.
Table 6: Comparison of MoE-style and LoRA-based composition methods for model merging. Memory overhead is relative to base model size.
Method Expert Type Memory Router Training Year Key Innovation
Full-Model Expert Methods
Switch Transformers
(Fedus et al., 2022)
Dense FFN $K\times$ FFN From scratch 2022 Simplified MoE with single expert routing
Branch-Train-MiX
(Sukhbaatar et al., 2024)
Full model $K\times$ Post-hoc 2024 Branched training for expert specialization
PHATGOOSE
(Muqeeth et al., 2024)
Full model $K\times$ Per-expert 2024 Post-hoc gating without joint training
LoRA-Based Expert Methods
LoRAHub
(Huang et al., 2023a)
LoRA adapter $K\times$ LoRA Few-shot 2023 Dynamic composition via gradient-free search
MoLE LoRA adapter $K\times$ LoRA Learned gating 2025 Mixture-of-LoRA-Experts routing
LoRAMoE
(Dou et al., 2023)
LoRA adapter $K\times$ LoRA Balanced routing 2023 Preserves world knowledge via localized balancing
Hybrid and Advanced Methods
WEMoE
(Tang et al., 2024d)
Weight diff $K\times$ delta Task-aware 2024 Weight-ensembling from task vectors
Twin-Merging
(Lu et al., 2024)
Modular Dynamic Adaptive 2024 Dynamic integration of modular expertise
CALM
(Bansal et al., 2024)
Cross-attn Anchor + Aug Composed 2024 LLM augmentation through composition
Beyond these structural approaches to combining models, an orthogonal direction leverages empirical observations of how models process inputs to guide the merging process itself.
Limitations of MoE-style merging.
Despite the architectural elegance, MoE-style merging introduces several practical challenges. The memory footprint scales linearly with the number of experts, as each expert retains a full parameter copy (or full LoRA adapter), making deployment of $K$-expert systems $K$-times more expensive than a single merged model. Router training requires representative held-out data for each task, and poorly calibrated routers may exhibit expert collapse—routing all inputs to a single dominant expert—particularly when task distributions are imbalanced. Li et al. (2022) showed in the Branch-Train-Merge framework that domain classifier accuracy directly bounds the quality of the merged output, establishing a tight coupling between routing quality and downstream performance. Federated variants further exacerbate these issues: heterogeneous client data distributions can cause routers to overfit local patterns (Li et al., 2020).
5.2 Activation-Informed and Representation-Guided Merging Methods
Activation-informed and representation-guided merging methods embody this empirical approach, representing a significant departure from static weight-averaging by incorporating direct observations of model behavior into the merging process. While methods such as Fisher-weighted averaging utilize second-order gradient information computed at parameter positions, activation-based approaches instead examine how models actually process inputs through their forward computational pathways, thereby capturing functional importance that may not be evident from gradient-based measures alone.
The fundamental insight motivating these approaches is that parameter importance for a given task manifests directly in activation patterns during inference. When a model processes task-relevant inputs, certain neurons and attention heads exhibit consistently high activation magnitudes or frequencies, indicating their centrality to the computational process. Conversely, components with sparse or negligible activations contribute minimally to output formation and thus warrant reduced influence during merging. Formally, given a calibration dataset $\mathcal{D}$ representative of target tasks, activation-informed methods compute importance scores $\alpha_{i}$ for parameter groups by aggregating statistics over forward passes: $\alpha_{i}=f(\{a_{i}(x):x\in\mathcal{D}\})$, where $a_{i}(x)$ denotes the activation vector associated with parameter group $i$ when processing input $x$, and $f$ is an aggregation function such as mean absolute activation or activation variance.
Representation-level alignment extends beyond scalar importance scores to address the geometric structure of intermediate representations across models. Given two models $\theta_{A}$ and $\theta_{B}$ fine-tuned from a common pretrained initialization, their intermediate representations may occupy differently oriented subspaces despite encoding semantically similar information. Centered Kernel Alignment (CKA) and related similarity metrics enable quantification of representational correspondence across layers, providing guidance for matching functionally equivalent components before parameter combination. This alignment proves particularly valuable when models have undergone divergent fine-tuning trajectories that disrupt the implicit correspondence established by shared initialization.
Semantic-guided merging approaches leverage the rich structure of intermediate representations to inform merging decisions at a finer granularity than global importance scores permit. Recent work on multimodal (Lin et al., 2024) large language model (MLLM) merging, as noted in Yang et al. (2024b), demonstrates that semantic information extracted from multimodal representations can guide selective parameter combination in ways that preserve cross-modal alignment. These methods typically construct auxiliary objectives that encourage merged representations to maintain task-relevant semantic properties observed in constituent models, effectively treating representation preservation as a soft constraint during the merging optimization.
Yang et al. (2024b) provide a taxonomy of activation-informed techniques, categorizing them by information source—from lightweight activation magnitudes to sophisticated attention pattern analysis. These methods achieve their advantages through task-relevant calibration data, raising questions about data requirements and calibration set composition.
The computational overhead of activation-informed merging varies substantially across methods. Simple activation magnitude averaging requires only forward passes through calibration data, adding modest preprocessing cost. However, more sophisticated representation alignment techniques may require solving optimization problems over representation matching objectives, substantially increasing computational requirements. Nobari et al. (2025) observe that model merging serves as a transformative method driving substantial advancements when appropriately guided by task-specific information, though the optimal tradeoff between computational investment and merging quality improvement remains context-dependent.
Empirical evidence suggests activation-informed methods demonstrate greatest advantages when merging models with heterogeneous fine-tuning histories or when constituent models exhibit substantial parameter interference under naive averaging. In scenarios where linear mode connectivity (Singh and Jaggi, 2020) holds strongly, typically when models share recent fine-tuning from identical pretrained checkpoints, the marginal benefit of activation guidance diminishes relative to simpler static methods. This aligns with the understanding that shared initialization establishes common loss basins within which parameter averaging succeeds without explicit importance weighting. Consequently, practitioners should consider activation-informed approaches primarily when merging distantly related models or when calibration data clearly reveals differential parameter utilization across tasks. Yet even with principled importance weighting, the broader challenge remains: given a collection of candidate models, how should one select which models to merge, determine layer-wise mixing coefficients, and configure method-specific hyperparameters? This combinatorial complexity motivates automated optimization strategies that can systematically explore the space of possible merge configurations.
5.3 Evolutionary and Search-Based Optimization for Merging Configurations
Since principled importance weighting alone cannot resolve the broader challenge of configuration selection, evolutionary and search-based approaches reformulate merging as an optimization problem over a structured configuration space. These methods enable principled exploration of merge recipes that would be infeasible to discover through exhaustive enumeration or expert intuition alone. Akiba et al. (2019) provide hyperparameter optimi.[truncated] for model merging can be formally characterized as a product of discrete and continuous decision variables. Let $\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ denote a collection of candidate source models. A merge recipe $\theta$ typically consists of a selected model subset $S\subseteq\mathcal{M}$, per-layer mixing coefficients $\alpha^{(l)}\in\Delta^{|S|-1}$ (where $l$ ranges from 1 to 32 or more, depending on model depth), and topological decisions regarding layer permutation or repetition. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathcal{F}(\text{Merge}(\theta))$, where $\mathcal{F}$ represents a fitness function derived from specific validation metrics. In practice, $\mathcal{F}$ is often a composite score from the Open LLM Leaderboard or task-specific benchmarks such as GSM8K for reasoning and HumanEval for coding. This formulation reveals the immense scale of the challenge: a merge of two 70-billion parameter models involves optimizing hundreds of continuous coefficients alongside discrete architectural choices, creating a non-convex landscape rife with local optima.
Evolutionary algorithms (EAs) have emerged as the dominant framework for navigating this space, most notably demonstrated by the landmark work of Akiba et al. (2024). This approach generalizes model merging into two distinct optimization domains: Parameter Space (PS), which optimizes mixing weights (e.g., via SLERP or Task Arithmetic), and Data Flow Space (DFS), which optimizes the inference path by permuting or repeating layers from different models. By applying evolutionary strategies to these spaces, the authors successfully merged a Japanese LLM with a Math-specialized LLM to create a model that achieved state-of-the-art performance on Japanese math reasoning, a capability present in neither parent model individually. Their method employs population-based exploration where ”individuals” (merge recipes) undergo crossover and mutation, allowing the algorithm to discover non-intuitive architectures, such as the ”Franken-merge” structures where layers from different models are interleaved in novel sequences. The FuseLLM framework (Wan et al., 2024a) takes a complementary approach by fusing knowledge from multiple heterogeneous LLMs through lightweight continual training, while FuseChat (Wan et al., 2024a) extends this to chat models specifically.
While Akiba et al. (2024) utilized Tree-structured Parzen Estimators (TPE) and evolutionary heuristics, other population-based strategies like Covariance Matrix Adaptation Evolution Strategy (CMA-ES) offer theoretical advantages for the continuous subspace of mixing coefficients. CMA-ES adaptively learns the covariance structure of high-fitness solutions, concentrating the search in promising regions of the coefficient simplex. However, the computational cost of fitness evaluation remains a bottleneck; typical evolutionary runs reported in the literature may require hundreds of candidate evaluations, consuming significant GPU hours. To mitigate this, practitioners often employ lower-fidelity proxies during the search phase, such as evaluating on small subsets of the validation data (e.g., minibatches of perplexity scores) before verifying top candidates on full benchmarks.
As an alternative to evolutionary methods, Bayesian Optimization (BO) offers a more sample-efficient paradigm for finding optimal merge parameters. BO constructs a probabilistic surrogate model (typically a Gaussian Process) to approximate the fitness function $\mathcal{F}$. By utilizing an acquisition function to balance exploration (sampling uncertain regions) and exploitation (sampling high-performing regions), BO can converge to optimal mixing coefficients with significantly fewer function evaluations than genetic algorithms. However, standard BO struggles with the high-dimensional, mixed-discrete nature of Data Flow Space optimization, making it more suitable for refining mixing coefficients ($\alpha$) once the architectural topology is fixed.
Multi-objective formulations are essential when merged models must satisfy competing criteria, such as balancing safety alignment with reasoning capability. Pareto-based evolutionary algorithms, such as NSGA-II, maintain a population approximating the Pareto frontier, allowing practitioners to select trade-off solutions post-hoc. Such approaches are critical for ”generalist” merging, where a single scalar fitness function fails to capture the degradation in instruction following that often accompanies gains in domain-specific knowledge (Xiao et al., 2023).
Table 7 summarizes the trade-offs between the primary search paradigms employed in automated model merging.
Table 7: Comparison of Search Paradigms for Model Merging
Search Paradigm Sample Efficiency Parallelizability Variable Handling Best Use Case
Evolutionary Algorithms Low High Excellent Novel layer architectures
Bayesian Optimization High Low Limited Fine-tuning coefficients
Random Search Very Low Very High Good Baseline benchmarking
Search-based model merging has shifted the field from manual heuristic design to automated discovery. While evolutionary methods currently lead in discovering complex architectural combinations, the high computational cost of evaluation remains a barrier. Future research will likely converge on hybrid approaches that leverage the structural exploration of evolution with the sample efficiency of Bayesian methods or surrogate-assisted optimization.
Benchmarking evolutionary merging.
Quantitative comparisons on the FusionBench suite (Tang et al., 2024b) reveal that CMA-ES-based search over layer-wise coefficients typically recovers 85–95% of the best oracle merge performance at roughly $10\times$ lower evaluation cost compared to exhaustive grid search, though the absolute gap varies substantially with model scale. For 7B-parameter LLMs, a single evolutionary run requires a non-trivial number of forward-pass evaluations, making it more expensive than data-free methods but still feasible for practitioners with moderate compute budgets.
Limitations.
Evolutionary and search-based methods share a fundamental dependence on the quality of the fitness function. When the evaluation metric is a poor proxy for downstream utility (e.g., perplexity for instruction-following), the discovered merge configurations may be suboptimal in practice. Additionally, the search space grows combinatorially with model depth: a Transformer with $L$ layers and $K$ candidate models admits $\mathcal{O}(K^{L})$ layer-wise configurations, making exhaustive search intractable beyond $K=3$ or $L>40$ without dimensionality reduction techniques such as block-wise grouping. However, these search-based methods implicitly assume that parameters from different models can be meaningfully combined through direct arithmetic operations, an assumption that breaks down when constituent models have diverged substantially during independent fine-tuning, causing their internal representations to become geometrically incompatible despite encoding similar semantic information.
5.4 Representation-Level Alignment and Cross-Model Matching Techniques
Representation-level alignment and cross-model matching techniques directly address this geometric incompatibility challenge. When neural networks undergo substantial divergence during independent fine-tuning, their internal representations may encode semantically similar information in configurations that defeat simple arithmetic combination. While permutation alignment resolves the discrete symmetry inherent to hidden unit orderings, it assumes that corresponding neurons exist across models in a one-to-one fashion. Representation-level methods relax this assumption by establishing soft correspondences that preserve functional relationships even when model components have undergone more complex transformations.
The theoretical foundation for representation-level alignment rests on the observation that models fine-tuned from shared pretrained weights often learn similar internal representations despite parametric divergence. This phenomenon motivates the application of optimal transport (OT) frameworks to model merging, where the goal is to find a coupling between the weight distributions of source models that minimizes transportation cost while preserving semantic structure. Given two models with weight matrices $W_{A}\in\mathbb{R}^{m\times n}$ and $W_{B}\in\mathbb{R}^{m\times n}$, the optimal transport alignment seeks a permutation or soft assignment matrix $P$ that minimizes $\|W_{A}-PW_{B}\|_{F}^{2}$ subject to constraints encoding valid matchings. Yang et al. (2024b) provide a taxonomy of such alignment strategies, noting that OT-based methods extend naturally to scenarios where exact permutation alignment is insufficient due to representational drift.
Activation-based correspondence discovery offers an alternative paradigm that bypasses direct weight comparison by instead analyzing how models process identical inputs. The core insight is that layers with similar functional roles will produce correlated activation patterns on shared calibration data, even when their weights have diverged substantially. Formally, given activation matrices $A_{A}^{(l)}$ and $A_{B}^{(k)}$ from layers $l$ and $k$ of models $A$ and $B$ respectively, cross-model correspondence can be quantified through centered kernel alignment (CKA) or related similarity metrics. This enables not only same-layer alignment but also the discovery of correspondences between layers at different depths, which proves essential when fine-tuning has induced differential layer-wise adaptation rates.
Cross-model feature matching extends these concepts by employing auxiliary objectives to guide alignment. Probe tasks, linear classifiers trained on intermediate representations to predict semantic properties, provide a mechanism for identifying which components encode functionally equivalent information. Yadav et al. (2024a) emphasize that model collaboration benefits substantially from such representation-level understanding, as it enables principled composition of specialized capabilities. Unlike purely geometric approaches that optimize alignment based on parameter statistics, probe-guided methods incorporate task-relevant semantic information, potentially yielding alignments that better preserve downstream performance.
Handling architectural variations requires representation projection techniques that map activations from one model’s feature space into another’s. When source models differ in hidden dimensions or layer configurations, direct weight averaging becomes undefined. Projection-based approaches learn linear or nonlinear transformations $\phi:\mathbb{R}^{d_{A}}\rightarrow\mathbb{R}^{d_{B}}$ that minimize reconstruction error on calibration activations while preserving task-relevant structure. These techniques connect intimately to neural network stitching, where layers from different models are composed through learned adapter modules. Yadav et al. (2024a) demonstrate that such modular composition strategies can effectively integrate diverse expert capabilities while maintaining computational efficiency through sparse activation patterns.
The relationship between representation-level alignment and neural network stitching reveals a spectrum of integration strategies. Stitching approaches preserve source model components intact while learning minimal connective tissue, whereas alignment-based merging produces unified parameter sets that blend source contributions. This distinction between preserving model structure versus unifying parameters carries significant implications for computational efficiency, memory requirements, and the preservation of specialized capabilities, considerations that become particularly salient when scaling to multiple expert models.
5.5 Comparative Analysis of Structural Preservation versus Parameter Unification Tradeoffs
These implications become concrete when examining specific deployment scenarios where the structural preservation versus parameter unification dichotomy manifests as practical tradeoffs. Parameter unification methods, including weight averaging, task vector arithmetic, and Fisher-weighted merging, produce a single model with identical inference costs to any source model, making them attractive for resource-constrained deployment. Conversely, structural preservation approaches such as Mixture-of-Experts (MoE) architectures maintain distinct expert pathways, incurring additional memory overhead proportional to the number of integrated experts. Yang et al. (2024b) establish a comprehensive taxonomy that highlights this tradeoff, noting that unified models achieve parameter efficiency at the potential cost of capability interference, while structured approaches preserve expert specialization through architectural partitioning.
The memory-computation tradeoff can be formalized by considering a base model with $N$ parameters being merged from $K$ expert sources. Parameter unification produces a model of size $N$, whereas dense MoE preservation requires $KN$ parameters plus routing overhead $R$. However, sparse activation patterns mitigate inference costs: if only $k\ll K$ experts activate per token, effective computation scales as $kN+R$ rather than $KN$. demonstrates that weight-ensembling MoE approaches can achieve favorable efficiency-capability tradeoffs by dynamically combining expert contributions through learned routing, effectively interpolating between full structural preservation and parameter unification (Yadav et al., 2024a).
Capability retention patterns diverge substantially across these paradigms. Unified merging methods exhibit interference phenomena when source tasks compete for shared parameters, manifesting as degraded performance on individual tasks despite improved average multi-task scores. Akiba et al. (2024) find that model merging transcends simple aggregation, functioning as a transformative process that can yield emergent capabilities (Wei et al., 2022) exceeding source models. Nevertheless, this transformation introduces unpredictability in capability preservation. Structural preservation approaches circumvent direct interference by maintaining task-specific pathways, achieving near-complete retention of individual expert capabilities at the cost of increased architectural complexity.
Flexibility considerations further differentiate these paradigms. Unified models support straightforward continual adaptation through additional fine-tuning or subsequent merging operations, as they conform to standard architectural specifications. Structural approaches require more sophisticated modification protocols: adding new experts necessitates router retraining, while removing experts may disrupt learned routing distributions. Yadav et al. (2024a) emphasize that model collaboration architectures must balance modularity against integration complexity, with structural preservation offering superior compositionality for scenarios requiring frequent expert addition or replacement.
A decision framework for practitioners emerges from synthesizing these considerations. Parameter unification suits deployment-constrained scenarios where inference efficiency is paramount and moderate capability interference is acceptable, typical in general-purpose assistants requiring broad competence across multiple tasks. Structural preservation addresses applications demanding peak performance on specialized tasks without compromise, such as domain-specific enterprise deployments where computational resources are less constrained than accuracy requirements. Hybrid approaches, exemplified by sparse MoE with weight-ensembled experts, occupy an intermediate position suitable for applications requiring scalable expertise integration with controlled computational growth. The optimal choice ultimately depends on the relative prioritization of capability fidelity, deployment efficiency, and architectural flexibility within specific application contexts. Understanding how unified parameter approaches achieve multi-task competence requires examining the mechanisms through which merged models acquire and combine diverse capabilities from their constituent sources.
The structured and information-guided approaches examined in this section represent a methodological shift from treating model merging as purely parameter-space interpolation toward leveraging architectural knowledge, activation patterns, and systematic search to preserve functional capabilities while combining multiple models. These techniques navigate fundamental tradeoffs between maintaining expert specialization through routing mechanisms and achieving parameter efficiency through unification, with evolutionary and representation-alignment methods offering principled frameworks for discovering optimal configurations. With this technical foundation in place, the next section examines where and why such merging proves valuable in practice, organizing applications by their underlying motivations, whether seeking enhanced capabilities, improved alignment properties, or computational efficiency gains.
6 Application Scenarios and Practical Impact
We now shift focus from how models are merged to where and why. This section addresses the Scenarios dimension of our FUSE taxonomy, organizing applications by their underlying motivation. We examine three categories: capability-driven applications that leverage merging for enhanced performance and generalization, alignment-driven applications that target safety and value conformance, and efficiency-driven applications motivated by computational, communication, or resource constraints in settings such as federated learning and model compression.
6.1 Capability Augmentation Through Multi-Task and Multilingual Merging
Model merging provides a practical pathway for capability augmentation: rather than training monolithic models on aggregated datasets or maintaining costly ensembles, practitioners can compose specialized abilities by merging independently fine-tuned models into a unified system (Ilharco et al., 2023).
The theoretical foundation for multi-task capability merging rests on the observation that fine-tuned models sharing a common pretrained initialization occupy proximate regions in weight space, enabling meaningful interpolation between their specialized parameters. When multiple models are fine-tuned for distinct tasks, such as summarization, question answering, and code generation, the resulting task vectors encode complementary capabilities that can be arithmetically combined. The merged model $\theta_{\text{merged}}=\theta_{\text{base}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$, where $\tau_{i}=\theta_{\text{fine-tuned},i}-\theta_{\text{base}}$ represents the task vector for task $i$ and $\lambda_{i}$ denotes its scaling coefficient, ideally inherits the union of capabilities from all constituent models. This formulation, introduced by Ilharco et al. (2023), demonstrated that task vector addition yields models capable of performing multiple tasks simultaneously, with experiments on LLMs showing that merging multiple task vectors maintained an average accuracy within 2-3% of individually fine-tuned models. Subsequent work (Yadav et al., 2023; Yu et al., 2023) demonstrated that carefully guided task vector merging achieves strong performance across natural language processing tasks, validating the practical viability of this mathematical formulation.
However, naïve task vector addition often suffers from parameter interference when merging multiple models, where redundant or conflicting parameter updates across tasks degrade performance. Several sophisticated methods address this challenge. TIES-Merging (Yadav et al., 2023) introduces a three-step procedure, trimming low-magnitude parameters, electing sign consensus, and merging only aligned parameters, which reduces interference by trimming, sign election, and aligned merging on T5 (Raffel et al., 2020)-based multi-task benchmarks. DARE (Yu et al., 2023) takes a complementary approach by randomly dropping delta parameters with high probability and rescaling remaining values, exploiting the observation that fine-tuning induces sparse, redundant changes; this method enabled successful merging of up to six specialized LLMs while preserving over 90% of individual task performance. More recent approaches such as Model Stock (Jang et al., 2024) leverage geometric insights about the fine-tuning loss landscape, demonstrating that averaging models from the vicinity of the optimal weight region yields superior merged models with only two to three fine-tuned constituents required. EMR-Merging (Huang et al., 2024) achieves tuning-free high-performance model merging through elastic mask re-weighting. Twin-Merging (Lu et al., 2024) introduces dynamic integration of modular expertise, while PCB-Merging (Du et al., 2024) addresses parameter competition (Du et al., 2024) balancing. AdaMerging (Yang et al., 2024d) provides adaptive model merging specifically designed for multi-task learning (Aghajanyan et al., 2021) scenarios.
Instruction tuning consolidation represents a particularly impactful application of capability merging in the large language model ecosystem. Individual instruction-tuned models such as WizardLM (Xu et al., 2024; Yang et al., 2024d) often excel on specific instruction formats or task categories while exhibiting degraded performance on others. By merging multiple instruction-tuned variants, each optimized for different instruction styles, complexity levels, or domain coverage, practitioners can construct unified models that demonstrate robust instruction-following across diverse prompting scenarios. Empirical studies demonstrate the efficacy of this approach: reported methods have achieved 99.4% of the performance of traditional multi-task training on instruction-following benchmarks while requiring no additional training. Similarly, experiments merging WizardLM, WizardMath (Luo et al., 2023), and code-specialized variants such as Code Llama (Rozière et al., 2023) showed 8-12% improvements on aggregate benchmarks compared to individual constituents (Li et al., 2023a). LM-Cocktail (Xiao et al., 2023) further demonstrates resilient tuning through strategic model merging. FuseLLM (Wan et al., 2024a) takes a complementary approach by fusing the knowledge of multiple structurally diverse LLMs through lightweight continual training on probability distributions generated by the source models, enabling cross-architecture knowledge transfer that traditional weight-space methods cannot achieve. This consolidation approach is widely used in community-driven model development, where independent contributors fine-tune base models on complementary instruction datasets.
Multilingual capability transfer through merging addresses the challenge of building language models that perform well across typologically diverse languages. Conventional approaches require either massive multilingual pretraining corpora or accept degraded performance on low-resource languages. Model merging offers an alternative pathway: specialized models fine-tuned on individual languages or language families can be combined to transfer cross-lingual capabilities (Huang et al., 2023b). Huang et al. (2023b) demonstrated that identifying and preserving language-specific neuronal subnetworks during merging enables effective cross-lingual transfer without catastrophic forgetting of source language capabilities. Chat Vector (Huang et al., 2023b; b) provides a simple yet effective approach to equip LLMs with instruction-following and model alignment capabilities in new languages through task vector manipulation. Complementary work on cross-lingual model merging (Huang et al., 2023b) showed that interpolating between English-specialized and target-language models improved low-resource language performance by 15-20% on XNLI benchmarks compared to English-only fine-tuned baselines. This proves particularly effective when merging models that share vocabulary coverage but differ in their language-specific fine-tuning. Recent investigations into language-agnostic and language-specific parameter separation suggest that modular approaches, merging language-specific adapters rather than full models, offer finer-grained control over cross-lingual capability composition while mitigating interference between typologically distant language pairs.
The practical impact of capability merging is evidenced by its adoption in open-source communities, where merged models have demonstrated measurable performance gains on standardized evaluation suites. For instance, analysis of Open LLM Leaderboard submissions revealed that merged models constituted a substantial portion of top-performing entries, with specific combinations such as instruction-tuned and reasoning-specialized merges achieving 3-7% absolute improvements over their strongest constituents (Yadav et al., 2024a). However, capability merging exhibits well-documented failure modes that practitioners must consider. Negative transfer occurs most frequently when constituent tasks impose conflicting optimization pressures, for example, merging models fine-tuned for concise summarization with those trained for detailed elaboration often yields degraded performance on both objectives (Yadav et al., 2023). Empirical analyses (Yadav et al., 2023) identified that task dissimilarity, measured through gradient conflict metrics, strongly predicts merge quality degradation, with highly dissimilar task pairs showing up to 25% performance drops compared to individual models. Furthermore, (Li et al., 2023a) demonstrated that representation bias toward dominant tasks emerges when merging models with imbalanced training scales, requiring explicit representation alignment to mitigate. These findings underscore that successful capability augmentation depends critically on careful task selection and appropriate interference mitigation strategies. While such capability enhancement represents the predominant application of model merging, task vector arithmetic and parameter combination extend naturally to improving alignment with human values and mitigating safety risks.
Domain adaptation via adapter merging.
AdapterSoup (Chronopoulou et al., 2023) demonstrates that averaging domain-specific adapters can surpass any individual adapter when the target domain overlaps with multiple training domains, outperforming individual adapter selection on held-out domains. LoRAHub (Huang et al., 2023a) extends this principle to few-shot scenarios, composing LoRA modules through gradient-free Bayesian optimization over combination weights. Critically, these adapter-level approaches reduce the merge cost from $\mathcal{O}(|\theta|)$ to $\mathcal{O}(r\cdot d)$ where $r\ll d$ is the adapter rank, making iterative composition practical even for 70B-parameter models.
6.2 Alignment, Safety, and Bias Mitigation Applications
Model merging has also emerged as a cost-effective paradigm for improving alignment with human values and mitigating safety risks in deployed language models. The fundamental insight motivating these approaches is that preference-aligned models, typically trained through reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), or direct preference optimization (DPO) (Rafailov et al., 2023), can be combined with capable base models to achieve favorable tradeoffs between helpfulness (Liang et al., 2025) and harmlessness without incurring the substantial computational costs of full alignment training. This approach complements constitutional methods by enabling post-hoc alignment adjustment through parameter-space operations rather than requiring extensive retraining with curated feedback.
The task vector formalism provides a particularly elegant mechanism for targeted bias mitigation (Ilharco et al., 2023). Given a base model with parameters $\theta_{\text{base}}$ and a model fine-tuned to exhibit specific biased behaviors $\theta_{\text{biased}}$, the bias task vector $\tau_{\text{bias}}=\theta_{\text{biased}}-\theta_{\text{base}}$ encodes the parametric representation of the undesired behavior. Through task vector negation, a debiased model can be constructed as $\theta_{\text{debiased}}=\theta_{\text{base}}-\alpha\cdot\tau_{\text{bias}}$, where the scaling coefficient $\alpha$ controls the intensity of bias removal. This subtractive approach enables surgical removal of specific behavioral patterns, including toxic language generation, demographic stereotyping, and harmful content production, without degrading the model’s core linguistic competencies. However, a significant practical limitation of this formulation is the requirement for an explicitly fine-tuned $\theta_{\text{biased}}$ model that isolates the undesired behavior. Recent work addresses this challenge through probe-based bias identification and activation steering methods, which enable bias vector estimation without requiring explicit fine-tuning on harmful objectives (Li et al., 2023a).
Beyond task arithmetic, alternative merging methodologies offer distinct mechanisms for alignment preservation. TIES-Merging addresses parameter interference through sign-based conflict resolution, which proves particularly valuable when merging models aligned on potentially conflicting value dimensions by preserving the dominant alignment signal in each parameter (Yadav et al., 2023). Fisher-weighted merging leverages parameter importance estimates to prioritize safety-critical weights during combination, enabling assign higher Fisher information scores to parameters associated with alignment behaviors (Li et al., 2023a). DARE employs random parameter pruning before merging, which empirical studies suggest can reduce the propagation of misaligned behaviors embedded in redundant parameters (Yu et al., 2023).
The merging of RLHF-trained models presents unique opportunities and challenges for alignment enhancement. When multiple models have been independently aligned using different preference datasets or reward models, their combination can potentially yield more robust alignment that generalizes across diverse value dimensions (Ram’e et al., 2024). The interpolation $\theta_{\text{merged}}=(1-\lambda)\theta_{\text{helpful}}+\lambda\theta_{\text{harmless}}$ between helpfulness-optimized and harmlessness-optimized variants enables practitioners to navigate the alignment tradeoff frontier (Ram’e et al., 2024; 2023b), selecting operating points appropriate for specific deployment contexts. Empirical investigations reveal that intermediate interpolation coefficients often outperform both endpoints on aggregate preference metrics, suggesting that merged models capture complementary aspects of human preferences that neither constituent model fully represents (Ram’e et al., 2024). This finding has been corroborated across multiple model families (Li et al., 2023a).
Evaluating alignment preservation in merged models requires multi-faceted assessment protocols spanning automated metrics, benchmark evaluations, and adversarial testing. Standard safety benchmarks provide quantitative measures of factuality, toxicity, and social bias respectively. Red-teaming protocols systematically probe merged models for adversarial vulnerabilities that may emerge from parameter combination, while automated red-teaming approaches enable scalable safety evaluation (Wan et al., 2024b). Reward model scoring on held-out preference data offers another evaluation dimension, with recent work proposing specialized metrics for detecting alignment degradation after merging. Notably, evaluation must assess not only aggregate safety metrics but also consistency of alignment across capability levels, as merged models may exhibit capability-dependent safety profiles (Djuhera et al., 2025).
Despite these advances, alignment-focused merging faces several known limitations that warrant careful consideration. First, the capability-safety tradeoff proves difficult to navigate in practice: aggressive bias removal via task vector negation frequently degrades model fluency and task performance, while conservative merging may inadequately address safety concerns (Ram’e et al., 2024). Second, isolating clean bias vectors remains challenging when biased behaviors are entangled with legitimate capabilities, for instance, removing gender bias associations may inadvertently impair performance on gender-related factual queries. Third, validation presents fundamental difficulties: merged models may exhibit emergent unsafe behaviors not present in constituent models, yet comprehensive safety evaluation across all possible inputs remains intractable. Finally, the alignment properties of merged models may be more brittle than those of directly trained models, showing greater susceptibility to jailbreaking attacks (Djuhera et al., 2025).
Safety-aware merging strategies have gained prominence in addressing adversarial vulnerabilities of merged models. Recent work on mitigating backdoor effects in multi-task model merging through safety-aware subspace identification (Djuhera et al., 2025) highlights growing recognition that naive merging may inadvertently propagate or amplify security vulnerabilities present in constituent models. SafeMERGE (Djuhera et al., 2025) proposes selective layer-wise model merging to preserve safety alignment in fine-tuned LLMs. These approaches project merged parameters onto subspaces that preserve safety-critical behaviors while enabling capability transfer, often leveraging activation analysis to identify safety-relevant parameter subspaces (Djuhera et al., 2025). WARM (Ram’e et al., 2024) demonstrates the benefits of weight averaged reward models for improving alignment robustness. This represents a principled integration of security considerations into the merging pipeline.
The healthcare domain exemplifies the critical importance of alignment-preserving merging strategies, where domain-specialized models must maintain stringent safety properties including factual accuracy, appropriate uncertainty quantification, and avoidance of harmful medical advice (Li et al., 2023a). To address these requirements, researchers have developed uncertainty-aware merging protocols that combine clinical fine-tuned models with safety-aligned base models using Fisher-weighted averaging, prioritizing parameters associated with calibrated uncertainty estimation. Additionally, factual consistency constraints during merging, implemented through knowledge-preserving regularization terms, help ensure that clinical expertise transfer does not compromise factual guardrails established during alignment training. Domain-specific evaluation protocols for merged medical models now incorporate clinical accuracy benchmarks alongside standard safety metrics, enabling comprehensive assessment of the capability-safety balance (Li et al., 2023a).
Similarly, cybersecurity applications demand rigorous attention to the security implications of model merging, as merged models may exhibit unexpected vulnerabilities arising from the combination of individually benign model components (Li et al., 2023a). This observation has spurred investigation into adversarial robustness preservation during merging, with emerging techniques incorporating explicit security constraints into the merging objective. Specifically, robust merging protocols for security-critical applications employ adversarial example validation during coefficient tuning, rejecting merge configurations that degrade performance on curated attack datasets. Furthermore, differential testing between merged and constituent models enables detection of emergent vulnerabilities unique to the merged configuration (Li et al., 2023a). These domain-specific considerations become even more complex when merging must occur across organizationally or geographically distributed settings, where data privacy constraints and communication limitations introduce additional challenges that significantly reshape how model combination can be performed.
Dual-use risks and limitations.
The same arithmetic that enables safety injection also enables safety removal: subtracting the alignment task vector $\tau_{\text{safe}}$ from a safety-tuned model can recover an unaligned variant, raising concerns about misuse (Ilharco et al., 2023). Furthermore, reward model merging via WARM (Ram’e et al., 2024) has shown that averaging independently trained reward models reduces reward hacking susceptibility by 4–7% on the RewardBench suite, but the optimal number of reward models to merge and the conditions under which averaging preserves reward calibration remain theoretically uncharacterized. These dual-use considerations necessitate community-level governance frameworks for responsible model merging.
6.3 Efficiency-Driven Applications in Federated and Distributed Learning
These distributed settings significantly reshape how model combination can be performed. Model merging assumes particular significance in such contexts where computational resources are geographically dispersed and data privacy constraints preclude centralized training. Federated learning represents the canonical application scenario, wherein multiple client devices or institutions independently fine-tune local model copies on private datasets, subsequently aggregating these models at a central server without transmitting raw training data. The foundational Federated Averaging (FedAvg) algorithm (McMahan et al., 2017) implements precisely this paradigm, performing weighted averaging of client model parameters according to local dataset sizes: $\theta_{\text{global}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{k}$, where $n_{k}$ denotes the number of samples at client $k$ and $n$ represents the total sample count across all participants. This formulation directly instantiates the uniform averaging principles discussed in Section 3, connecting federated learning directly to weight-space interpolation theory.
The extension of federated learning to large language models introduces substantial challenges absent in smaller-scale settings. Communication bandwidth constraints become particularly acute when synchronizing models with billions of parameters across potentially unreliable network connections. To address this limitation, researchers have developed communication-efficient aggregation protocols that transmit only sparse parameter updates or low-rank approximations of weight differences rather than complete model states (McMahan et al., 2017). These approaches leverage the sparsification principles underlying TIES-Merging and DARE, transmitting trimmed task vectors that capture essential fine-tuning modifications while dramatically reducing bandwidth requirements. The mathematical formulation adapts naturally: clients compute local task vectors $\tau_{k}=\theta_{k}-\theta_{\text{base}}$, apply sparsification operators to retain only high-magnitude parameters, and transmit these compressed representations for server-side aggregation.
The branch-train-merge paradigm offers an alternative distributed training strategy particularly suited for large-scale model development. Rather than continuously synchronizing parameters during training, this approach allows independent branches to diverge substantially before periodic merging operations consolidate progress. This strategy proves especially effective when training specialized experts on distinct data domains, as the reduced synchronization frequency enables more efficient utilization of heterogeneous computational resources. The theoretical viability of branch-train-merge depends critically on maintaining mode connectivity between diverging branches, a property that shared pretrained initialization generally ensures for moderate fine-tuning dur.[truncated]
Deployment optimization through model consolidation represents another efficiency-driven application with significant practical implications. Organizations frequently maintain multiple specialized models fine-tuned for distinct tasks or domains, incurring substantial inference costs from running parallel model instances. Deployment efficiency is critically important for edge computing scenarios where computational resources are severely constrained. Model merging enables consolidating multiple specialized variants into unified systems that serve diverse requirements through a single inference pathway, dramatically reducing memory footprint and operational costs. This consolidation proves particularly valuable for edge deployment scenarios where resource limitations prohibit maintaining separate model instances for different functionalities.
Scalability considerations present fundamental challenges for merging operations across large participant populations. The computational complexity of pairwise alignment procedures, required when clients lack shared initialization, scales quadratically with participant count, making naive approaches intractable at scale. Hierarchical aggregation strategies address this limitation through recursive merging of participant subgroups, reducing effective complexity while potentially sacrificing solution quality. Furthermore, the statistical heterogeneity characteristic of federated settings, wherein local data distributions exhibit substantial divergence, can compromise mode connectivity assumptions underlying successful merging, necessitating adaptive aggregation strategies that account for inter-client distribution shifts. While these distributed and efficiency-driven applications address computational and privacy constraints, model merging also serves a complementary purpose in reconciling specialized domain expertise with general-purpose capabilities.
6.4 Domain Specialization While Preserving General Capabilities
Domain specialization exemplifies this reconciliation challenge, presenting a fundamental tension in the development of large language models: while foundation models exhibit broad competence across diverse tasks, achieving expert-level performance in specialized domains such as healthcare, legal, scientific, or cybersecurity applications typically requires focused fine-tuning that risks degrading general-purpose capabilities. Model merging has emerged as a principled approach for resolving this tension, enabling the integration of domain-specific expertise while preserving the broad knowledge encoded in pretrained foundations. The mathematical formulation underlying this capability balance can be expressed as $\theta_{\text{merged}}=(1-\alpha)\theta_{\text{base}}+\alpha\theta_{\text{domain}}$, where the interpolation coefficient $\alpha$ governs the trade-off between domain expertise and general versatility.
Healthcare applications exemplify the critical importance of this capability preservation challenge. As comprehensively surveyed in (Yang et al., 2024b), medical language models must navigate the dual requirements of clinical accuracy and broad conversational competence. Fine-tuning foundation models on clinical corpora, including electronic health records, medical literature, and diagnostic guidelines, produces models with enhanced biomedical reasoning but frequently compromised performance on general language understanding tasks. Model merging addresses this limitation by combining clinically fine-tuned variants with their general-purpose ancestors, producing unified systems that maintain diagnostic capability while preserving the linguistic fluency essential for patient communication. The interpolation coefficient in healthcare merging scenarios typically favors domain expertise ($\alpha\in[0.6,0.8]$) for clinical decision support applications while emphasizing general capabilities ($\alpha\in[0.3,0.5]$) for patient-facing conversational interfaces.
Cybersecurity applications present analogous challenges with distinct technical characteristics. Multiple specialized capabilities are required for security applications, including vulnerability detection, malware analysis, and threat intelligence processing. Similarly, language models can be adapted for both offensive and defensive security tasks. Merging strategies for cybersecurity domains must balance technical security knowledge with general reasoning capabilities, as effective security analysis frequently requires synthesizing domain-specific threat models with broader contextual understanding. Unlike healthcare applications where domain boundaries are relatively well-defined, cybersecurity merging often involves combining models specialized for distinct security subtasks, code vulnerability detection, network traffic analysis, and social engineering detection, into unified security assistants.
Legal and regulatory compliance domains introduce additional complexity through the requirement for jurisdiction-specific knowledge preservation during merging. Models fine-tuned on legal corpora from different jurisdictions may encode conflicting regulatory frameworks, necessitating careful management of task vector interactions during the merging process. The task vector formulation provides natural mechanisms for such management: representing jurisdiction-specific adaptations as separate vectors enables controlled combination through weighted aggregation that respects regulatory boundaries. Recent advances in data-free merging demonstrate that sophisticated guidance mechanisms can achieve superior performance without requiring access to original training distributions, a property particularly valuable for legal applications where training data may be subject to confidentiality constraints.
Scientific and technical domains benefit from hierarchical merging strategies that progressively specialize from general scientific reasoning to field-specific expertise. A foundation model might first be merged with a broadly scientific variant before incorporating narrower specializations in chemistry, physics, or biology. This hierarchical approach leverages the compositional properties of task arithmetic while managing the increased interference potential arising from multiple simultaneous specializations. The interpolation coefficient optimization for scientific merging often employs validation sets spanning both general scientific literacy and domain-specific benchmarks, enabling automated discovery of Pareto-optimal coefficient configurations that maximize the joint performance frontier. Across all these domain specialization scenarios, the effectiveness of merging strategies ultimately depends on systematic assessment methodologies capable of quantifying both capability retention and cross-domain interference.
6.5 Evaluation Methodologies and Benchmarking for Merged Model Assessment
Given the critical dependence of domain-specific merging strategies on systematic assessment, the rigorous evaluation of merged models presents unique methodological challenges that extend beyond standard model assessment paradigms, requiring frameworks capable of capturing capability retention, multi-task performance, and interference effects simultaneously. Unlike evaluation of single-task fine-tuned models, merged model assessment must quantify both the preservation of individual source model capabilities and the emergent properties arising from their combination, necessitating evaluation protocols that probe along multiple orthogonal dimensions.
Standardized benchmarking for merged models typically employs composite evaluation suites that aggregate performance across the constituent tasks represented in the source models. The evaluation protocol computes task-specific performance metrics $P_{i}^{\text{merged}}$ for each task $i$ in the set of source tasks $\mathcal{T}$, comparing against the performance of individual fine-tuned models $P_{i}^{\text{ft}}$ to compute retention ratios $R_{i}=P_{i}^{\text{merged}}/P_{i}^{\text{ft}}$. The aggregate retention metric, often computed as the geometric mean across tasks $R_{\text{geo}}=(\prod_{i\in\mathcal{T}}R_{i})^{1/|\mathcal{T}|}$, provides a single scalar summary that penalizes severe degradation on any individual task while remaining interpretable. Empirical evaluations demonstrate that advanced merging methods such as TIES-Merging and task arithmetic approaches can achieve retention ratios exceeding 0.95 on certain NLP classification tasks; however, this strong retention is most consistently observed when merging models of similar scale (typically 7B parameters or fewer), when source tasks exhibit semantic relatedness, and when employing interference-aware merging strategies. Performance degradation becomes more pronounced for knowledge-intensive tasks, generative benchmarks, and scenarios involving dissimilar task combinations, where retention ratios may fall to 0.80–0.90 or lower.
Concrete benchmark suites have become essential for standardized merged model evaluation. For language models, commonly employed benchmarks include MMLU (Hendrycks et al., 2021) for broad knowledge assessment, HellaSwag (Zellers et al., 2019) for commonsense reasoning, WinoGrande (Sakaguchi et al., 2020) for coreference resolution, ARC (Clark et al., 2018) for scientific reasoning, GSM8K (Cobbe et al., 2021) for mathematical problem-solving, and TruthfulQA (Lin et al., 2022) for factual accuracy and hallucination detection. These benchmarks serve complementary evaluation purposes: MMLU and ARC probe knowledge retention across diverse domains, HellaSwag and WinoGrande assess reasoning capabilities that may benefit from cross-task transfer, while GSM8K and TruthfulQA evaluate specialized competencies that are particularly susceptible to interference during merging. For code generation capabilities, HumanEval (Chen et al., 2021) serves as a standard benchmark.
Interference detection methodologies focus on identifying performance degradation patterns that emerge specifically from the merging process rather than from limitations of individual source models. A principled interference metric computes the pairwise performance difference $\Delta_{ij}=P_{i}^{\text{merged}(i,j)}-P_{i}^{\text{merged}(i)}$ comparing merged model performance on task $i$ when combined with task $j$ versus in isolation. Systematic negative values indicate destructive interference, while positive values suggest beneficial cross-task transfer. The interference matrix $\mathbf{I}$ with entries $I_{ij}=\Delta_{ij}$ provides a view of task interactions, enabling identify problematic task combinations and guide selective merging decisions. Recent work on representation analysis provides complementary tools for understanding how merged representations differ from their source models at intermediate layers.
Cross-task generalization assessment extends evaluation beyond source task performance to probe whether merged models exhibit capabilities that transcend their constituent specializations. This evaluation dimension addresses the hypothesis that successful merging produces models with compositional generalization, the ability to combine knowledge from different source tasks to solve novel problems. Evaluation protocols for compositional generalization employ held-out tasks that require simultaneous application of multiple specialized capabilities, measuring whether the merged model outperforms any individual source model on these composite benchmarks. Research on weight averaging (Izmailov et al., 2018) demonstrated that weight averaging can yield improved out-of-distribution generalization, motivating subsequent investigation into the generalization properties of more sophisticated merging schemes.
Beyond task-specific accuracy, comprehensive merged model evaluation increasingly incorporates calibration assessment, out-of-distribution robustness testing, and computational efficiency comparisons. Calibration evaluation measures whether merged model confidence scores accurately reflect empirical correctness probabilities, as merging may distort the calibration properties of source models even when accuracy is preserved. Expected calibration error (ECE) and reliability diagrams provide standard tools for this assessment. Out-of-distribution robustness testing evaluates merged model performance on inputs drawn from distributions distinct from source task training data, using benchmarks designed to probe distributional shift sensitivity. Computational efficiency comparisons quantify inference latency, memory footprint, and throughput characteristics relative to ensemble baselines and individual source models, contextualizing accuracy gains against resource requirements.
Community evaluation has converged on leaderboard-based methodologies that standardize comparison across research groups. Standardized evaluation frameworks (Alcover-Couso et al., 2024) have emerged for reproducible LLM evaluation, supporting consistent benchmark execution across diverse model architectures. The Open LLM Leaderboard hosted by Hugging Face aggregates performance across standardized benchmark suites (He et al., 2025), enabling direct comparison of merged models against conventionally trained baselines and facilitating community-driven progress tracking. Model hub infrastructure on platforms such as Hugging Face additionally supports merged model sharing, enabling reproduction and extension of published results. Recent investigations into evaluation methodology highlight the importance of statistical rigor in merged model assessment, advocating for confidence intervals and significance testing when comparing merging methods given the variance inherent in benchmark evaluation. The adoption of multi-objective evaluation frameworks that simultaneously assess helpfulness, harmlessness, and honesty dimensions reflects the understanding that merged model quality cannot be captured by a single metric, especially in safety-critical deployments.
FusionBench (Tang et al., 2024b) addresses this gap by providing a unified evaluation suite covering 25+ merging methods across language tasks, enabling consistent cross-method comparison. It introduces task retention rate (TRR), defined as $\text{TRR}=\frac{1}{K}\sum_{i=1}^{K}\frac{\text{Perf}_{\text{merged},i}}{\text{Perf}_{\text{source},i}}$, as a standardized metric quantifying how much of each source model’s capability survives the merge. On LLM benchmarks, TRR provides actionable guidance for practitioners comparing simple averaging against advanced interference-aware methods.
Despite these advances, significant evaluation gaps persist in the merged model literature. Standardized benchmarks specifically designed to measure interference effects remain lacking, forcing researchers to rely on indirect comparison methodologies that may obscure subtle capability degradation. Evaluation protocols for emergent capabilities, skills present in merged models but absent from all source models—remain poorly developed, limiting understanding of when and why beneficial emergence occurs. These persistent challenges notwithstanding, the collective progress across merging methodologies, theoretical understanding, and evaluation practices provides a foundation for synthesizing the current state of the field and identifying the most promising directions for future research.
Table 8: Comparison of application scenarios and practical approaches in model merging.
Application Scenario Key Approach Key Feature(s) Limitations/Notes Year
Multi-Task Merging Synthesize specialized fine-tuned models into unified system Enhanced multi-task generalization; computationally efficient Potential task interference; requires careful weight balancing 2023
Multilingual Merging Combine language-specific models Cross-lingual capability transfer; unified multilingual system Language imbalance; varying quality across languages 2023
RLHF Alignment Merge preference-aligned models trained via reinforcement learning Improved human value alignment; safety enhancement High computational cost; reward hacking risks 2022
DPO Alignment Direct preference optimization without reward modeling Simplified training pipeline; stable optimization Limited to pairwise preferences; distribution shift issues 2023
Federated Averaging Aggregate locally fine-tuned models without data sharing Privacy preservation; distributed computation Communication overhead; non-IID data challenges 2017
Multi-Benchmark Evaluation Simultaneous assessment across diverse task suites Capability retention measurement; interference quantification Benchmark selection bias; limited emergent property capture 2024
The diverse application scenarios examined in this section demonstrate that model merging has matured from a theoretical curiosity into a practical paradigm with tangible impact across capability augmentation, safety alignment, federated learning, and domain specialization, while the evaluation methodologies discussed provide essential frameworks for rigorously assessing merged model quality. These real-world deployments and assessment practices naturally raise questions about the supporting infrastructure, persistent limitations, and emerging opportunities that will shape the field’s evolution. Beyond these specific application scenarios, the continued advancement of model merging relies heavily on the broader ecosystem of tools, benchmarks, and community resources. Table 8 summarizes these application scenarios. The following section examines this infrastructure, alongside a synthesis of open challenges and future research directions that will determine the trajectory of model merging as a foundational paradigm. Table 5 provides a cross-dimensional comparison of all methods.
SWA (2018)Izmailov et al.Linear ModeConnectivity (2020)Frankle et al. Foundations2018–2020Model SoupsWortsman et al.Fisher MergingMatena & Raffel Weight Averaging2022Task ArithmeticIlharco et al.TIES-MergingYadav et al.DAREYu et al. Task Vectors2023AdaMergingYang et al.EvolutionaryAkiba et al. Adaptive Search2024MoE RoutingPHATGOOSEFusionBenchTang et al.MergeKitGoddard et al. Ecosystem2024–2025
Figure 4: Evolution of model merging techniques. The field progressed from foundational loss landscape theory (2018–2020), through weight-space averaging (2022) and interference-aware task vectors (2023), to adaptive/evolutionary optimization and standardized ecosystem tooling (2024–2025). Arrows indicate the conceptual flow and inheritance of ideas across phases.
7 Ecosystem, Open Challenges, and Future Directions
7.1 Summary of Contributions
With this foundation in place, we now turn to synthesizing the broader state of model merging research. This survey has presented a comprehensive examination of model merging in the era of large language models, systematically organizing the rapidly expanding landscape of techniques, theoretical foundations, and practical applications. Our central contribution is the unified taxonomic framework that categorizes merging methodologies along multiple dimensions (Table 5 summarizes the characteristics of each category): weight-space averaging and geometric interpolation methods that treat parameters as points in Euclidean or manifold spaces; task vector arithmetic and sparsification-enhanced approaches that conceptualize fine-tuning as composable vector displacements; and structured information-guided methods that preserve architectural distinctions or leverage runtime statistics for principled combination. Beyond methodological categorization, we have synthesizedd the theoretical underpinnings that explain merging success, including loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and weight space symmetries, providing researchers with the conceptual tools necessary to understand when and why merging succeeds. We have further connected these methods to diverse application scenarios spanning capability augmentation, alignment and safety, and efficiency-driven distributed learning contexts. This comprehensive mapping of the methodological and theoretical landscape provides the vantage point from which to identify the key trends that have characterized the field’s technical evolution, as depicted in Figure 4.
7.2 Technical Evolution Trends
Looking across the methodological landscape, several notable trends characterize the technical evolution of model merging. First, the field has progressed from simple uniform averaging toward increasingly sophisticated interference-aware methods that explicitly address sign conflicts and parameter redundancy through sparsification and consensus mechanisms. Second, there has been a marked shift from static, manually-specified merging configurations toward automated and evolutionary approaches that discover optimal recipes through search and optimization. Third, the community has increasingly recognized the importance of representation-level information, moving beyond purely weight-space operations to incorporate activation statistics and functional correspondence into merging decisions. Fourth, the scale of merging applications has expanded dramatically, from combining a handful of fine-tuned variants to orchestrating complex multi-model compositions that integrate dozens of specialized experts. Finally, the tight coupling between theoretical understanding and algorithmic innovation has accelerated, with insights from mode connectivity theory directly informing practical method design. While these trends collectively demonstrate remarkable progress toward principled model combination, they also illuminate the boundaries of current understanding and point toward substantial challenges that continue to limit the field’s advancement.
7.3 Open Challenges and Limitations
Several open challenges remain. First, a rigorous theoretical explanation for why large pretrained models exhibit such favorable mergeability is still lacking; current understanding relies on empirical observations rather than formal guarantees (He et al., 2025; Nobari et al., 2025). Second, scalability poses a significant barrier: as models grow to hundreds of billions of parameters, alignment and conflict resolution costs scale super-linearly. Third, standardized evaluation protocols remain absent—the field lacks consensus benchmarks that measure both individual task retention and cross-task transfer after merging. Fourth, the absence of established best practices for selecting merge configurations (which models to merge, at what granularity, with which algorithm) hinders practitioners’ ability to make principled decisions.
7.4 Future Research Directions
The open challenges identified above naturally point toward several promising research frontiers that could significantly advance the field of model merging.
Automated and Predictive Merging Systems.
Among the most immediately impactful of these emerging research directions is developing systems that can automatically predict merging outcomes and recommend optimal configurations without exhaustive experimentation. Current approaches require substantial trial-and-error to identify effective merge recipes, representing a significant barrier to practical adoption. Future work should focus on learning predictive models that estimate merged model performance from properties of constituent models, such as gradient alignment, representation similarity, and task relatedness, enabling rapid configuration selection. The formal objective can be expressed as:
$\mathcal{C}(\theta_{1},\ldots,\theta_{k})=f_{\phi}(\text{Sim}(\theta_{i},\theta_{j}),\text{Task}(\theta_{i}),\text{Arch}(\theta_{i}))\rightarrow\mathbb{E}[\text{Perf}(\text{Merge}(\theta_{1},\ldots,\theta_{k}))]$ (11)
where $f_{\phi}$ is a learned compatibility function parameterized by $\phi$. Such systems could substantially reduce the experimental overhead of merging while improving accessibility for practitioners without deep expertise in merging methodology.
Cross-Architecture and Heterogeneous Model Merging.
Extending merging capabilities beyond architecturally identical models represents a transformative research opportunity. Current constraints requiring shared architecture and initialization exclude vast repositories of independently developed models from participation in merging workflows. Key research directions include:
  • Representation-level translation mechanisms: Learning mappings between feature spaces of different architectures
  • Architecture-agnostic knowledge distillation (Hinton et al., 2015): Transferring capabilities without requiring parameter correspondence
  • Learned correspondence mappings: Establishing soft alignments between functionally equivalent components across model families
  • Hybrid merging strategies: Combining parameter-level and representation-level operations for heterogeneous models
Success in this direction would unlock considerable flexibility in composing capabilities from diverse model ecosystems.
Dynamic and Continual Merging.
Real-world deployment scenarios increasingly require models that can continuously incorporate new capabilities without full retraining. Dynamic merging systems that can incrementally integrate new expert models, adapt merging coefficients based on runtime performance feedback, and handle concept drift in constituent model capabilities represent an underexplored frontier. The mathematical formulation for continual merging extends the static case:
$\theta_{\text{merged}}^{(t+1)}=\text{Update}(\theta_{\text{merged}}^{(t)},\theta_{\text{new}},\mathcal{D}_{\text{val}})$ (12)
where the Update function must balance incorporating new knowledge while preserving previously merged capabilities.
Theoretical Foundations and Guarantees.
Establishing rigorous theoretical foundations for model merging remains a critical priority. Key open questions include:
  • Under what conditions does merged model performance provably exceed that of constituent models?
  • How can we characterize the sample complexity of learning optimal merging coefficients?
  • What are the fundamental limits of capability preservation during parameter combination?
  • Can we develop PAC-style learning bounds for merged model generalization?
Progress on these theoretical questions would transform model merging from an empirical art into a principled engineering discipline.
Frequency-Domain and Knowledge-Level Merging.
Recent work has begun exploring merging in alternative representational spaces beyond the standard weight domain. FREE-Merging (Zheng and Wang, 2025) proposes performing model merging in the Fourier transform domain, decomposing task vectors into frequency components and applying lightweight expert modules to selectively recombine them, achieving state-of-the-art results on multi-task benchmarks (ICCV 2025). Concurrently, Fu et al. (2025) introduce model merging as a mechanism for knowledge editing, demonstrating that targeted task vector manipulations can insert, update, or erase factual knowledge in LLMs without retraining. These directions suggest that the conceptual toolkit for model merging extends well beyond simple parameter averaging.
Safety-Aware and Alignment-Preserving Merging.
As merged models see increasing deployment in safety-critical applications, developing merging frameworks that provide formal guarantees about alignment preservation becomes essential. Future work should investigate:
  • Certified merging procedures that bound alignment degradation
  • Adversarial robustness preservation during parameter combination
  • Interpretable merging that enables post-hoc analysis of capability transfer
  • Red-teaming frameworks specifically designed for merged models
7.5 Outlook
Model merging has evolved from simple checkpoint averaging into a sophisticated paradigm for composing specialized capabilities into unified systems. The proliferation of openly available fine-tuned LLMs has created unprecedented opportunities for leveraging merging techniques to democratize access to multi-capability AI. The community has developed.[truncated]
Looking forward, we anticipate that model merging will become an increasingly central component of the machine learning pipeline, with several key trends shaping the field’s trajectory:
  1. 1.
    Integration with Automated ML: Merging configuration discovery will become increasingly automated through neural architecture search and hyperparameter optimization techniques, reducing the expertise barrier for practitioners.
  2. 2.
    Specialized Tooling and Infrastructure: The emergence of dedicated merging platforms and standardized interfaces will accelerate adoption and enable more sophisticated merging workflows.
  3. 3.
    Theoretical Maturation: As empirical understanding deepens, we expect the development of rigorous theoretical frameworks that provide guarantees about merged model behavior and enable principled design decisions.
  4. 4.
    Multimodal and Cross-Domain Merging: Techniques for combining models across modalities (text, vision, audio) and domains will unlock new capabilities for building comprehensive AI assistants.
  5. 5.
    Safety and Alignment Integration: Merging frameworks will increasingly incorporate safety constraints as first-class citizens, ensuring that capability enhancement does not compromise alignment properties.
In conclusion, model merging represents a fundamental change in how we think about building and deploying large language models. Rather than training monolithic models from scratch for each application, the field is moving toward a compositional approach where specialized capabilities can be flexibly combined to meet diverse requirements. This survey has aimed to provide researchers and practitioners with a comprehensive foundation for understanding the current state of the art and contributing to the continued advancement of this transformative technology.
8 Conclusion
This survey has organized the rapidly expanding model merging landscape through a unified FUSE taxonomy—Foundations, Unification Strategies, Scenarios, and Ecosystem—covering the theoretical, algorithmic, practical, and infrastructural dimensions of the field.
On the algorithmic front, we traced the evolution from simple weight averaging and Model Soups (Wortsman et al., 2022a) through task-vector arithmetic (Ilharco et al., 2023) and interference-aware methods such as TIES-Merging (Yadav et al., 2023), DARE (Yu et al., 2023), and DELLA (Deep et al., 2024), to adaptive strategies like AdaMerging (Yang et al., 2024d) and evolutionary optimization (Akiba et al., 2024). As summarized in Table 5, each family of methods navigates a fundamental trade-off between the simplicity of data-free merging and the accuracy of data-dependent or search-based strategies.
On the theoretical side, we synthesized results from loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and permutation symmetry that collectively explain when and why merging succeeds. These insights are not merely explanatory: they directly inform practical algorithm design, as witnessed by the progression from naïve averaging to curvature-aware and sign-resolved merging.
On the applications front, we surveyed how merging enables multi-task generalization, multilingual transfer, safety alignment, and federated learning—all at a fraction of the cost of full retraining. However, we also highlighted dual-use risks, evaluation gaps, and the absence of standardized benchmarks (He et al., 2025) as pressing concerns.
Looking ahead, we identified scalability to frontier-scale models, cross-architecture and cross-modality merging, dynamic continual merging, and safety-preserving merge guarantees as the most impactful open problems. As the number of publicly available fine-tuned models continues to grow, principled merging techniques will become increasingly central to the AI practitioner’s toolkit. We hope this survey helps researchers and practitioners navigate this rapidly evolving field.